A Complete Bibliography of Publications in Theory of Probability and its Applications: 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

31 May 2024 Version 3.01

Title word cross-reference

(α, q) [Sza12, Sza13]. 2 [Pin15]. 6 − j [BM16d]. A [Yak10, Yak15]. α [Mal18a]. α, 0 < α < 2 [EK19]. β ∈ (0, 0.5] [VT14]. B [ACS17]. χ^{2} [Ma17, Zac19]. d [Zai19]. δ [BP16]. GI|G|∞ [AR13]. I [VK18]. ∞ [Aib17]. K [AM19a]. L_{+}^{0} [Kra16]. L^{f} [CJJ11]. L^{p} [Jaj10]. L_{2} [Roz19b]. M [Lin18, Vol11, WHV18]. a [Fel10]. R^{+} [Maz12]. Z^{4} [HTV12]. R [YYZ13]. M ≪ N [GH11]. N [Mal18a, Prz15]. N/M → 0 [Bao15]. p ≥ 2 [BN19]. Q [Rao18, Rok19, Sza12, Sza13]. R [Shu17]. R_{Φ} [Gap11]. ρ [WH19]. s [CJJ11]. S_{w} [Gap11]. t [Sav15, Shv15]. U [Vol11, Yan12, Yan13]. V [BV16]. Z [WH14]. Z^{1} [PP17]. Z^{d} [PR19].

-Adic [Fel10]. -Branching [Mal18a]. -Demimartingales [Prz15]. -Dependent [Vol11]. -Differenced [AM19a]. -Distance [Zac19]. -Distribution [Sav15, Shv15]. -Estimators [Lin18]. -Free [ACS17]. -Function [VK18]. -Independence [Rao18]. -Invariant [YYZ13]. -Learning [Rok19]. -Mapping [Yak15]. -Mixing [WH19]. -NOD [WHV18]. -Norm [Roz19b]. -Optimal [BP16]. -Ornstein [Sza13, Sza12]. -Permutations [Yak10]. -Recurrent [Shu17]. -Risk [Zai19]. -Selfdecomposable [CJJ11]. -Smooth [Pin15]. -Spaces [Jaj10]. -Stable [Mal18a]. -Statistics [BV16, Vol11, Yan12, Yan13]. -Weakly [Maz12]. -Wiener [Sza12, Sza13].

100th [PSM14].

2014 [Yas15]. 2nd [Ali18c, Ali18a].

3rd [SP19].

60th [Kha13c].

70th [Kha14a]. 75th [Kha10].

80th [Kha12c, Kha13d, Mol12, Poz13, Vat11a]. 85th [Kha17b, Shi18b].

A. [Kha10, Kha12c, Kha13d, Kha16b, Kha17b, Mol12]. Abel [Yas15]. Abelian [Fel17, Fel18, Fel19, Yak19]. Absence [KP12, KP13, Nag12, Nag15a]. Absolute [KS10, Nag10, Yam12]. Abstracts [Ali18a, SPTS17, SP19, Ali18c]. Academician [Shi18b]. Accumulation [Kas16, PP17]. Accuracy [Nov14, Sak10, She13, She14b]. Achieving [Lyu12]. Action [Bak15]. Active [Zhu14]. Additional [WF18]. Additive [ZK18]. Adic [Fel10]. Advanced [SU11]. Aeppli [Ruk10]. Affine [Ser11]. Aggregate [BNB18]. Aleksander [Kha14a]. Aleksei [Kha15]. Algorithm [CPC13]. Algorithms [HY16]. Almost [Var19, ZN11]. Alternatives [NN12]. among [Ma13, Pin18]. Amounts [BNB18]. Analogues [Sen10b]. Analysis [AR13, BNB18, CK19, Dor10, KL11, Pal18, SU11, SW11, Tyu11b, WF18]. Analytic [ISF17a, Kuz11a]. Analytical [ZG17]. Anatolii [Ibr12]. Angular [CO11]. Annuities [ZG17]. Anomalous [Pal19]. Any [Lyk18]. Application [AP18, Pal18, She18b]. Applications [AR13, BNB18, BL12, Bor14, HM12, HM13, Lin18, MM13, NR16, Pin19a, Sak15, WHV18, WH19, Yam12]. Approach [DP13, Har16, LR10, SW11, Sta15, ZP18]. Appropriate [BK12, BK13a]. Approximate [Kud19]. Approximating [BL12]. Approximation [Bor15, GH11, HZ11, ISF19a, Osm10, Sen10a, Sen11, Sen15, Ser10, She13, She14b, Var15, VUC13, VP10, ZP18, ZG17]. Approximations [GS12, GH14, KV13, WF18, ZKKS17]. Arak [GEZ18]. Arbitrary [Bor17b]. Areas [VK16]. Array [Bor10]. Arrays [Wu15, Yon12]. Arrow [Kra16]. ASCLTs [ADES19]. Asian [NAKL17]. Asian-Type [NAKL17]. Asset [CKT16]. Assignments [CLS15]. Associated [Bak10, DP18b, Dem14, ISF19b, Nez11]. Assumption [Shk11b]. Assumptions [Roz18b]. Asymmetric [CS17]. Asymmetry [Kuz11c]. Asymptotic [Aly17, Bao15, BK17, BV16, DM13, GIKP15, KL11, Lin18, NK13, Osm10, Pal18, PG14, Pol18, Rag10, Sen12, Sen13, SS13, Sen16, Sha15, SCN12, SCN13, Syu10a, Syu14, TPP12]. Asymptotically [DSW19, PGU18, She11, WD16, Zai19]. Asymptotics [AA18, HZ11, KS15, Leb10, Lot15, NP16, NP13, Pus11, Roz19b, Sak15, WD16, Yak11, Yar18, vDZP10]. Asynchronous [VL13]. Attaining [Afa10, Afa11b]. Autoregression [Bol13, Bol19]. Availability [BK17]. Average [Kar13, Kar14]. Averages [Sid10, Tyu10]. Averaging [Khi14a, Sha15]. Award [CDD15, Kha16a].

B [Yas16]. B. [Bin14, Kha14b, PSM14, VZ16]. Backward [Bah10, Gla16, HO16]. Balking [Bel12]. Ball [NP16]. Banach [Fat14, Mam12, Mam13, Mam17, Pin15]. Band [JSS16]. Banded [BL12]. Bandit [Sav19]. Bang [YYZ13]. Barrier [AKG16, ZX19]. Barycenters [PST17]. Based [BN19, NSB10, PP10, SP18]. Basket [NAKL17]. Bayesian [Lis19, ZS13a]. Be [GU16]. Becker [Mas12]. Behavior [Aly17, BK17, GIKP15, Pal18, PG14, Rag10].

Behavioral [RRV16]. Belong [Syu14]. Benchmarks [YYZ13]. Benes [KL14]. Bergström [Syu10a, Syu10b]. Bergström-Chebyshev [Syu10a]. Bernoulli [Che11, Chi16, IMV15, NBS13, Zai15]. Bernstein [Fel12, GGW14]. Bernstein-type [GGW14]. Berry [ADES19, DM13, KS12, KS10, She11, WH19]. Besov [Rad10]. Beta [VP10]. Beta-Distribution [VP10]. between [BN19, BKS18, Iva15, Khi14b, Rok19, Sav17, Zac19, ZKK12, ZKK13]. Bibliography [Kho13]. Bifractional [Ma13]. Binomial [HZ11, Ma17, NC12, She18b, VUC13, WN11, Yak11, Zac19, ZS13b]. Binomial- [Ma17]. Binomial-Type [WN11]. Birth [vDZP10]. Birth-Death [vDZP10]. Birthday [Ali18b, Kha10, Kha12c, Kha13d, Kha13c, Kha14a, Kha17b, Mol12, Poz13, PSM14, Shi18b, Vat11a]. Blockwise [Nez11]. Boolean [Ser11]. Bootstrap [GH11]. Boris [Kov14]. Borovkov [Kha12c, Kha13d, Kha17b]. Bougerol [Vak12]. Bound [Kar14, KS10, Leb19, NC12, Pin19b, WH19, Yas10]. Boundaries [DSW19, Kaj17, Rag10]. Boundary [Afa19, BRU10, Bor17b, ISF14, ISF15, ISF17b, ISF18, RA13, Rai13, WD16]. Bounded [GM19, ISF17b, ISF18, Prz15, Var15]. Bounds [ADES19, BC10, Kar13, KD18, NAKL17, Pin19a, PGU18, Rok10, Rok12, Sak10, WF18, Yan12, Yan13, ZKKS17, Zha12, vDZP10]. Boxes [IMV15]. Branching [Afa10, AGGR15, Bul11, Bul15, Che15, Dya12, HTV12, KY19, Koz10, Leb10, Leb13, LV19, Mal18a, PR19, VW10, Vat11c, VL13, Vat15, Vat16, VD18a, VD18b, WH14, Yar11, Yar18]. Bridge [Lis19]. Bridges [BBE17]. Brings [Khi14a]. Bronius [Kha15]. Brownian [Afa11a, AB16b, AL19, BS11, BBE17, GP17, GU16, Lis19, Ma13, Mur11, NP13, NKL14, Sau19, Shi17a, Sin12, Spi12, Vak12, ZM13, ZS14]. BSDEs [NR16]. Bucy [BBD11].

Can [GU16, Roz18a]. Canonical [BV16, BB19, KT14, Vol11, WF18]. Capacity [Hol18]. Carlo [Kud19, Moh14]. Case [Bor18b, GM19, Kuz11c, Lin16, NKL14, PST17, Sta15, VD18b, BCG13, VD18a]. Cases [Osm10]. Catalytic [Bul11, Bul15]. Cauchy [CO11, Pla17, Sha12b]. Causes [PF12]. Censored [Top10]. Central [Bao15, Bul17, KS18, PG14, RMC14, Sen15, BS11, KP12, KP13, MS13, NS13, Sen11, Sen12, Sen13, SS13]. Certain [Mak17, MR15, NN11]. Chain [CPC13, Lyu12, PS10c, RA13, Shu14]. Chains [BC10, But14, ZKKS17]. Chance [AKG16]. Change [Asg14, Dar12, Dar14, Pol18, SS17]. Change-Point [Asg14, Dar12, Dar14, Pol18]. Changepoint [TPP12]. Channel [Hol18]. Channels [KH14, KH15]. Chaos [ZP18]. Characteristic [PGU18, Zha12]. Characteristics [Khi14a, Sav17]. Characterization [BP15, BCG13, Fel18, Fel19, Gus11, Hol17, LG11]. Characterizations [Rao18]. Chebyshev [BM12a, Sen10b, Syu10a, Syu10b]. Chebyshev-Type [BM12a]. Chernoff [ZM13]. Chervonenkis [Kha15]. Chung [PL10]. Chung-Type [PL10]. Claim [BNB18]. Claims [Kha13a, Rok12, ZX19]. Class [CJJ11, EST11, SZK13, SCN12, SCN13]. Classes [Gap11, GNT15]. Classical [Hol18, Sok14]. Classification [Bul15, Gla16]. Claude [Kha16a]. Clearing [KME18]. Close [Rod19]. Closed [GP10]. Closer [Khi14a]. CLT [BX16, Roz19a, She18a, She18b]. Clusters [Gol14]. Coalescing [Kon11a]. Coding [KH14, KH15]. Coefficient [Kry19].

Coefficients [GZ12, GZ13, SF11]. Coherent [Kon10, Kon11b]. Collector [DP18a]. Colored [AGGR15]. Colorings [Sem19]. Common [Roz16, Roz18b, Roz18a]. Communications [Bor19]. Commutant [Mal17]. Compact [Fel18, Gol13]. Compactness [HS16, Kha18a]. Compensated [TE16]. Compensated-Covariation [TE16]. Compensator [Gus18]. Complete [Bul15, KMM15, ZS13b]. Completeness [GS19]. Complex [GZ12, GZ13]. Complexity [Ser10]. Component [Tim19]. Components [Il’10, Yak15]. Compound [BM16b, BM16c, Bor18a, Bor19, She14b, She18a]. Computable [WF18]. Concentration [BS14, EZ13, GGW14, GEZ18, KD18, Tol17]. Condition [PST17, SL13]. Conditional [Afa14, Bul17, Kuz11b, Son16]. Conditions [AB12, AB13, AT19, Bah10, BX16, GNT18, Gri10, KV13, KD18, Nag11, Roz13, Sid10, Yam12]. Conference [Ali18a, Ali18c, Dor10, SPTS17, SP19]. Congratulations [Kha12a, Kha16b]. Connected [Sav15]. Connection [Iva15]. Consistency [GH11]. Constant [DM13, Gri19, KL11, KS10, ZX19]. Constants [She11]. Constructing [LB18]. Construction [Sen16]. Consumption [BP16]. consumptions [Kra16]. Containing [GD15]. Contingent [Kha13a, Rok12]. Continuity [CV10, Kru10, Yam12]. Continuous [DP13, LaG13a, Pol18, Var19, WN11, ZKKS17]. Continuous-Time [ZKKS17]. Contrast [RMC14]. Control [Bah10, HY16, Pal18]. Convergence [AKG16, BK12, BK13a, Deh13, FV10, Gap11, GT10, GP10, ISF15, IMV15, Kar17, KS11, KP12, KP13, Nag10, NS13, Roz13, Sak11, SZK13, She18a, She18b, Shu14, Su15, Tim18, Tyu11a, Var15, Var19, Yas10, Yas11, Yon12, vDZP10]. Converging [FKZ14, Lin16]. Convolution [Syu14].

Copulas [Sav15]. Corollaries [HS16]. Correlated [AL17b, Kas12, MR15]. Correlation [Suk10, WF18]. Cost [Sav19]. Costs [NR16]. Cosy [Lau11]. Countable [PST17]. Coupon [DP18a]. Covariance [BS11, Bao15, Dem14, GT10, ORS14, Yas18]. Covariances [SK11]. Covariation [TE16]. Cramér [BCG13, Koz14]. Cramér-Type [BCG13]. Criteria [Kha18a, RRV16, Sav17, Yar11]. Criterion [HS16, Irk11]. Critical [Afa10, Afa11b, GIKP15, LV19, VW10, VL13, VD18a, VD18b]. Cross [Bel17]. Crossing [Rag10, RA13, WD16]. Crossings [BN12, SK11]. Cubic [EST11]. Cumulative [Aly17]. Current [GS19]. Curved [WD16]. Customer [Gri19]. Customers [AB12, AB13]. CUSUM [Shi17a]. Cycle [Yak10, Yak18]. Cycles [Tim17]. Cylinders [MF15].

Dalang [Rok10]. Darmois [Ibr13]. Data [KPP14]. Death [vDZP10]. Debreu [Kra16]. Decayed [Kaj17]. Decomposable [Vat15, Vat16]. Decomposition [Shi19a]. Decompositions [PP10, Sau19]. Decorrelation [AM19b]. Decrease [Roz18b, Roz18a]. Decreases [Roz16]. Decreasing [Roz15, Zai15]. Dedicated [Hol16]. Defect [Zai15]. Defined [Kas16]. Definition [ISF17a]. Deformed [PN15]. Degenerate [Koz17, Sha12a]. Degrees [Shv15]. Delayed [ZX19]. Delocalization [GNT18]. Demimartingales [Prz15]. Densities [Ale10, Bor18b, Koz17, NSB10, Rok10]. Density [BN19, Khi10, Mal17, Mum11, Mum12, Sak15, Yak19]. Department [Shi18a]. Dependence [Gri10, Maz18, Ser10]. Dependent [GNT15, Leb19, Sha15, Tyu10, Vol11, Wu15]. Description [CM17]. Detection [BK13b, Dar14, Pol18, SS17, TPP12, YMW15]. Determinate [Lyk18]. Determined [Jaj10]. Developments [AL17a]. Deviation [Aib17, BM12b, BM13a, BM13b, BM14a, BM14b, BM16b, BM16c, Mum11, Mum12, NBS13, NP13, Roz10, Roz15, Roz18b, Roz18a, Shk11b]. Deviations [AB16a, AL17b, BN19, Bak15, BCS11, Bor10, Bor19, CL10, Koz10, Koz14, Pus11, Roz14, Roz16, Roz19b, Shk11a, Var15, Zan13, dRdC17]. Diagonally [KT14]. Difference [Fel17]. Differenced [AM19a]. Different [Lyu12]. Differentiability [Mor12]. Differentiable [Son16]. Differential [BK16, BE14, BK12, BK13a, Mam12, Mam13, Mam17, MZ18, Pal18, Sha12a]. Differentiation [GD15]. Diffusing [Kon11a]. Diffusion [Ark15, Bel17, BK12, BK13a, CL10, GD15, Har16, ISF17a, IS12, Mak17, Noa10, RRV16, Sak10, Sak11, Sta15, DP18b]. Diffusion-Type [CL10]. Diffusions [BKS18, Hol19, Koz17, MPR10, Pal19]. Dilations [Gre10]. Dimension [Ser10]. Dimensional [Ark15, BE14, Bul11, Hol19, ISF14, IMV15, Kuz11b, Mal17, Sta15, Vat15, WF18]. Discontinuous [Bor18b, Per12]. Discounting [Sav19]. Discrete [ADES19, AKG16, Fel19, Gre10, KPP14, Nov11, PR11, Rok12, ZP18]. Discrete-Time [Nov11]. Discrimination [Rod18, Rod19]. Disintegrations [LaG13a]. Disorder [Ali13, Shi17a, ZS13a, ZS14]. Dispersion [SCN12, SCN13]. Distance [AP18, Bur16, Ser11, Zac19, Zhu11]. Distances [BKS18, Sav17]. Distributed [Bar13, Bor10, GM15, SF11]. Distribution [AKG16, Aly17, BN19, Bar13, BN12, BD13, Bor15, Bor17b, Bys15, FV10, GZ12, GZ13, Har16, HZ11, Koz10, Kru14, Lyu12, Mal17, Maz18, Maz12, Mum11, Mum12, Mur11, NSB10, NC12, PMS19, Rod19, Roz16, Roz18b, Roz18a, Ruk10, Sak19, Sav15, Ser11, Shv15, Son16, VK18, VP10, ZS13b]. Distributions [BRKF11, BKS18, BV16, BB19, CJJ11, Fel19, Gam18, IMV15, Il’10, Kha18a, Khi10, Kor14, Kor17, Leb10, Moh14, Noa10, Nov14, Rao18, Rod18, She13, She14b, She18a, SCN12, SCN13, SL13, Syu14, Tim18, VT14, Vat15, WN11, Yak11, Yam12, YB11, Zac19, Zha12]. Disturbed [BK12, BK13a]. Dividend [ZX19]. Divisibility [JGM16]. Divisible [BRKF11, Yak11, Yam12]. Domain [AL19, GD15, ISF17b, ISF18, RMC14]. Doubly [Kor17]. Drift [ADES19, BE14, BN12, Bor15, Mur11, Sin12, ZM13]. Driven [MZ18, NR16, Rad15]. Dual [Gus11]. Duality [IS12]. Dynamical [Dor10]. Dynkin [Yas16]. Dyson [BBP18].

E. [Yas16]. Editors

[Fat15, Kha16c, Prz16, Shu12]. Effect [Zai15]. Efficiency [Kar13, Kar14]. Eigenvalue [Bao15, JSS16]. eigenvalues [BBP18]. Elasticity [KL11]. Elementary [Kry16]. Elements [KT14, Tyu10, Uly16]. Ellipticity [Sak19]. Embedded [GU16]. Embedding [JKV14]. Empirical [Bak15, BK17, Deh13, FV10, Mum11, Mum12, SP18]. Empty [IMV15]. Enough [Roz18a]. Ensembles [HS16]. Entries [GNT15, ORS14]. Entropy [BO13, Kuz11b]. Environment [Afa10, Afa13, Afa14, Afa17, Afa19, CL10, Dya12, Koz10, VW10, Vat11c, VD18a, VD18b]. Environments [VL13]. Epidemic [Mas12, Yas10, Zhu14]. Equation [BK12, BK13a, Mam12, Mam13, Pal18, Sha12b]. Equations [BK16, BE14, BKS18, But14, DP18b, Gla16, GS19, LR10, Mam17, MZ18, Pla17, Rad15, Rai13, Sha12c, Sha12a]. equilibria [Kra16]. Equivalent [FP10]. Erdos [BO13]. Ergodic [AKG16, ACS17, But14, Jaj10]. Ergodicity [Bel12, CPC13]. Erratum [Ali18c, VD18a]. Error [WF18]. Esseen [ADES19, DM13, KS12, KS10, She11, WH19]. Esseen-Katz [She11]. Estimate [Bul16, Dem14, KS11, Kry19, Zac19].

Estimates [Ale10, EZ13, Gap11, KP12, KP13, NSB10, NS13, Sen11, Sen13, Sen15, Sha15, She13, She18a, She18b, Shi19a, Sok14, Syu10b, Zai19]. Estimating [Osm10]. Estimation [AB16b, Bor18b, KPP14, RMC14, Sak10, Sta15, YMW15]. Estimator [ADES19, NBS13]. Estimators [BN19, Deh13, GL13, Lin16, Lin18, LB18, NK13, NKL14, Sak15]. Euclidean [Mal17]. Euler [MZ18]. Evolution [ISF19a, LR10, Pla17, Rad15]. Evolutions [Gre10]. Evolving [VL13]. Exact [NP13, Pin19a, Roz19b, She11, WD16]. Exactness [Osm10, Sen10a, Sen11, Sen15]. Exciting [Ali13]. Excursions [HKP16, KPH14, Pit18]. Existence [Kra16, PST17]. Exit [AL19, GD15, KP19]. Expansion [Bao15, Rad18, Sen12, Syu10b]. Expansions [BV16, Osm10, Sen13, SS13, Sen16, Syu10a, Syu14]. Expectation [Nag15a]. Expectations [ISF15, ISF19a]. Expected [Leb19, SF11, Vos17]. Explicit [BM16a, CM17, LB18]. Explosion [Aly17]. Exponential [BNB18, BK12, BB19, BM12a, Bys15, CV10, Iva15, KL14, Mak17, SV18, Vos17, Yar11, BK13a]. Exponentially [Kaj17, Roz15]. Exponentials [GM19, Shu17]. Extended [BNB18, BNP12, Var19]. Extension [KL14]. Extensions [Mal18b, VK18]. Extinction [Top10, VW10]. Extremal [Gol14, Sav17]. Extremes [Leb13, Zhd16, ZP18].

factorial [AP18]. Factorization [CJJ11, Fel12, Kud19, Kuz11a]. Faculty [Shi18a]. Families [HS16, ISF19b, TE16, VD17]. Family [GL13, Sav19]. Fast [Roz18a]. Faster [Roz16]. Fatou [FKZ14]. Federation [CDD15]. Field [Bul16, MS13, Sha16]. Fields [AM19a, Bak10, BS14, DP13, DP19, KS18, Ma17, Pit18, Pus11, Roz19b, RMC14, Sak15, Ser10, Sha15]. Filtered [ZS13a]. Filtrations [Jaj10, Kru10, Lau11]. Final [Har16, Mas12, Vat11c]. Financial [CKT16, KS16, KS17, KME18]. Finetti [MS19]. Finite [IMV15, Il’10, Kas16, Lyk18, Tim18, Vat15]. Finite-Dimensional [IMV15, Vat15]. First [AL19, Bor17b, DSW19, GD15, Kaj17]. First-Passage [DSW19]. Fisher [Lin16]. Fit [NSB10, SP18]. Fitting [Maz18]. Fixed [Lyu12]. Flow [AT14, AB16a, AT19, AA18, Kry19]. Flows [GS19]. Fluctuation [Cor11]. Fluctuations [JSS16, ORS14]. Fokker [BKS18, Sha12b]. Follower [Rok19]. Forms [BKNP15, SS13]. Formula [EK19, Khi10, Per12]. Formulation [BRU10]. Forward [Bah10, CM17]. Four [VK16]. Fourier [Rad18]. Fractal [RMC14]. Fractional [ADES19, AB16b, AL19, BCS11, CK19, DP18b, Ma13, NKL14]. Free [ACS17, BRU10]. Freedom [Shv15]. Frequencies [Ruk10]. Function [BK17, Boy12, FV10, Gus11, Mor12, Mum11, Mum12, NN11, Roz18a, Ser11, Son16, VK18, ZS13b]. Functional [KS18, Kas16, MS13, Sha16, Vat16]. Functionals [BRU10, BC10, GP17, GD15, ISF15, ISF19a, Lyu10, NP13, SV18, SP18]. Functions [Dem14, EZ13, Gol13, GEZ18, Kel10, KD18, Li15, Naz10, ORS14, Pal19, Per12, PGU18, Rad10, Ser11, Yak16, Zha12, Zor15]. Fundamental [CKT16]. Funding [NR16].

G [Aib17]. Galton

[Afa11b, Dya12, GIKP15, Nag15b, Top10]. Game [Rok19]. Gamma [YB11]. Gaps [Khi14b, ZKK12, ZKK13]. Gas [Lag13b]. Gaussian [AL17b, BKNP15, Bys15, Fat14, Fel12, Hol17, Hol18, HKP16, Il’10, KP19, KPH14, KT14, LaG13a, Maz18, MS13, Moh14, Naz10, Pit18, Roz19b, Sak15, Sha16, SK11, VC10, Zan13, Zhd16, ZP18, ZN11]. General [BX16, GL13, HO16, Ker16, Rad15, Shi18a].

Generalization [Bor17a, GP10, Sza12, Sza13]. Generalized [AKG16, BE14, BS18, CJJ11, HS16, JKV14, Khi10, Kor14, Mam12, Mam13, Roz13, Sak19, TPP12, Zac19]. Generated [GS19]. Generating [Yak16]. Generators [Hol19]. Genetic [CPC13]. Geometric [GU16, Koz10, Tyu11b]. Geometry [Zor15, Zor18]. Gittins [Sav19]. Given [Ali18a, Ali18c, Sav17, SP19, Tim19]. Global [MPR10, PF12, Sak11, She18a, She18b]. Gnedenko [Bin14, Ker16, Kov14, PSM14]. Goodness [NSB10, SP18]. Goodness-of-Fit [NSB10, SP18]. Governed [Pal19]. Graphs [Bur16, EK18, Zhu11]. Grigelionis [Kha15]. Group [PP10]. Groups [Fel17, Fel18, Fel19, Gol13, Shu17]. Growth [Yar11].

Half [PMS19]. Half-Line [PMS19]. Hamiltonian [LMM13]. Hamming [Ser11]. HARA [CM17]. Hardy [SL13]. Having [HKP16, KPH14, Roz18b]. Hawkes [EK18]. Heat [Shi19a]. Heavy [BD13, GZ12, GZ13, Maz18, Nov14, SW11, VT14]. Heavy-Tailed [Nov14, GZ12, GZ13]. Heavy-Traffic [SW11]. Hedging [IS12, Kha13a]. Hermite [AM19b, Sen10b]. Hermitian [BX16]. Heston [BM16a]. Heterogeneous [BNB18]. Heyde [Fel18]. hierarchical [BBP18]. High [AB12, AB13, Afa10, Afa11a, Afa11b, Afa13, Afa17, Sen16, Top10, WF18, Zhd16, ZP18]. High-Dimensional [WF18]. Higher [Ale10, KP12, KP13, Rod19]. Hilbert [NP16, Rad15]. Himself [Pro18]. Hitting [CO13, Vak12]. Holevo [Kha14a]. Homogeneity [BN19]. Homogeneous [Kru12, Kru13]. Hopf [Kud19, Kuz11a]. Hybrid [KH14, KH15]. Hyperbolic [CO13, Kor14]. Hypergraph [Sem19]. Hypersphere [HZ10]. Hypotheses [Bol13, LS19, Rod19, Sav17, ZM13].

Hypothesis [Bol19, NN12].

I-Cosy [Lau11]. I. [Poz13]. I.I.D. [BCG13]. Ibragimov [Poz13, Shi18b]. Identically [Bar13, GM15, SF11]. Identity [Kru14, Kuz11a, Vak12]. II [BM13b, BM16c, HKP16, ISF18, KH15, Mum12, Roz15, Shu11, Vat16]. III [BM14a, Shu13]. Imaging [Sak10, Sak11]. Immigration [GIKP15]. Impatient [AB12, AB13]. Improved [She13]. Improvement [KP12, KP13]. Improvements [Com18]. Improving [VP10]. Incomplete [RRV16]. Increasing [Kaj17]. Incremental [BNHS17]. Increments [BM14b, Khi10, Khi14a, Kru10, Kru12, Kru13, SV18]. Independence [BP15, Rao18]. Independent [BN19, BM14b, EZ13, Fel10, Fel17, GM15, ISF15, ISF19a, Kha18a, Khi10, Khi14a, Kru10, Kru12, Kru13, Kru14, Kuz11c, MF15, NN12, Roz14, Roz15, Roz16, Roz18b, Roz18a, Roz19a, SV18, SF11, She13]. Index [EK19, Sav19]. Indicator [Dem14]. Indices [Gol14]. Induced [KPZ18]. Inequalities [BS14, BB19, BM12a, Bys15, CLS15, GGW14, GEZ18, Kol13, Li15, Mak17, Nag15b, Pin15, Tol17, WHV18, ZS13b]. Inequality [AM19b, KS10, She11, She18b]. Inference [Nov14]. Infinite [Bor15, JGM16, Kon11a, Kuz11b]. Infinite-Dimensional [Kuz11b]. Infinitely [BRKF11, Yak11, Yam12]. Infinity [Yak11]. Influence [Bin14, LMM13]. Information [Asg14, CS17, Dar14, Shi18a, Vos17, WF18]. Inhomogeneous [ZKKS17]. Initial [ISF14, ISF15, ISF17b, ISF18, Lin16, Rai13]. Initial-Boundary [ISF14, ISF15, Rai13]. Input [AT14, AB16a, AA18, Kry19]. Insider [BZ17]. Insurance [BNB18]. Integer [Gam15]. Integrability [Shu11, Shu13]. Integral [BRU10, GP17, Kol12, LR10, NBS13, RA13]. Integrals [Bys15, KS18, Sha16].

Integrating [Bys15]. Integration [BOGP14, Kry16]. Integro [BK16, Bor10, Bor17a, Bor18a, Roz19a]. Integro-Differential [BK16]. Integro-Local [Bor10, Bor17a, Bor18a, Roz19a]. Intensity [BK13b, KY19]. Interaction [Bul16]. Interference [AKG16]. International [Ali18a, Ali18c, Dor10, SP19, SPTS17]. Interpretation [Bel17]. Interruptions [AT19]. Intervals [BBE17, DM13]. Intrinsic [WH14]. Introduction [IKS18]. Invariance [Afa10, Afa11b, Bor14]. Invariant [YYZ13]. Inverse [BL12]. Investment [BP16, BM16a, CS17, RRV16]. Irreducible [PP10, Shu14]. Isotropic [Pit18]. IT-Copulas [Sav15]. Itô [Per12]. Iterated [Prz15].

Japanese [SU11]. Joint [Gus18, PP17, Ruk10, YMW15]. Jointly [KMM15]. Jump [Afa11a]. Jumps [HO16].

Kac [NP16]. Kagi [Spi12]. Kalman [BBD11]. Katz [She11]. Kendall [Leb19]. Kernel [BN19, Ma13, NSB10, Sak15, Shi19a]. Kernel-Type [NSB10]. Key [VT14]. Khintchine [JGM16]. Kiefer [NP16]. Known [Leb19, Sza12, Sza13]. Kolmogorov [BKS18, Gla16, Sha12b]. Kullback [Bak15].

Lamperti [DP19]. Lamperti-Type [DP19]. Laplace [Fat14, Zha12]. Large [AB16a, Aib17, Bak15, BCS11, Bor10, BM12b, BM13a, BM13b, BM14a, BM14b, BM16b, BM16c, Bor19, Chi16, CKT16, Gap11, HKP16, KP19, Koz10, Koz14, KPH14, Sen10a, Shk11a, Shk11b, Var15, Yas10, Zan13, Zhu14, dRdC17]. Lattice [Bul11]. Law [Che11, Chi16, Fel12, Gap11, GNT18, Gus18, PL10, Prz15, Ruk10, Sen10a, Yas10, Zhu11, Zhu14, ZS13b]. Laws [BM16d, Gol13, Kor14, NN11, Syu14, Tyu10].

Leader [Rok19]. Leading [Che15]. Learning [Rok19]. Leftmost [Che15]. Lemma [FKZ14]. Lengths [Tim17, Yak10]. Letter [Fat15, Prz16]. Letters [Shu12]. Level [Afa10, Afa11b, Afa13, Afa17, MS13, Sha16, Top10]. Levels [Lyu12]. Lévy [CV10, Cor11, Gla16, ISF19b, Iva15, JGM16, Kaj17, Khi14a, Kud19, Ma13, Mal18b, MZ18, NN11, Vos17]. Lévy-Driven [MZ18]. Life [Bor12, Bor13, Bul18, Dor10, Kha12b, Kha13b, Kha17a, Kha18b, KMS14, Kov14, Man15, Mol12, Nik13, PS14, PS10a, PS10b, She14a, Shi11, Shi15b, Shi15a, Shi18a, Su13, Vat11b]. Life-Information [Shi18a]. Like [Shi19a]. Likelihood [MS19]. Limit [AB12, AB13, Afa14, Bak10, BS11, Bao15, Bor18a, BS18, Bul17, Che15, DP13, EST11, Gam15, GNT15, GM15, GM19, Gri10, ISF15, ISF17a, ISF17b, ISF18, KS18, KY19, Kon10, Kon11b, KP12, KP13, Kor14, Kor17, Kuz11c, MS13, Mum11, Mum12, NS13, Nez11, RA13, RMC14, Sav15, Sen11, Sen12, Sen13, SS13, Sen15, Ser11, Sha16, Shk11b, Shu11, Shu13, Tim18, Vat16, Vol11, Wu15, Yak10, Yan12, Yan13, ZK18, ZN11]. Limiting [NN11]. Line [BOGP14, PMS19]. Linear [Bao15, BX16, BP15, DP13, Fel10, Gol14, GL13, ISF19b, JSS16, KS15, LMM13, Mam17, MF15, Pal18, Sta15]. Liouville [Pla17]. Liptser [Shi19b]. Littlewood [GEZ18]. Load [AB12, AB13]. Local [Bor10, Bor17a, Bor18a, Deh13, Gam15, GNT18, Khi14a, Khi14b, KV13, Nag15a, PR11, Rag10, Roz19a, Sha15, ZKK12, ZKK13]. Locally [Fel18, MR15, Nov11, NN12]. Log [LG11]. Log-Normal [LG11]. Logarithm [PL10, Prz15]. Logarithmic [Iva15]. Long [SF11, VD17]. Long-Tailed [SF11]. Loss [BRU10]. Lower [Kar13, Kar14, Rok10, Sak10]. LR [WF18]. Lyapunov [KS11, Li15, Tyu11a, Tyu12, Tyu13].

M [Aib17]. M/G/1/ [Aib17]. Many [VD17, Zor15]. Mapping [Var19, Yak15]. Mappings [Tim19]. Marc [JS15]. Market [RRV16, Rok12]. Markets [BP16, CKT16, NR16]. Markov [BC10, BM16d, Bul16, But14, CPC13, GGW14, Gre10, Har16, Kas16, Kel10, Khi14b, Lyu12, Mas12, Nov11, PS10c, RA13, Sav19, Shu14, ZKKS17]. Markovian [Dya12, SZK13]. Marshall [BNB18]. Martingale [KL14, Kol12, PST17, PMS19, Rok10]. Martingales [NR16, Pin15, PR11, TE16]. Maruyama [Mag10]. Mass [Kol13]. Massive [HKP16, KPH14, Pit18]. Matching [AP18]. Matérn [Pus11]. Mathematical [HM12, HM13]. Mathematician [Yas15]. Mathematics [KS16, KS17, Shi18a]. Matrices [Bao15, BX16, BL12, GT10, GNT15, JSS16, ORS14, Yas18]. Matrix [Asg14, ORS14, Pal18, Shv15, Suk10]. Matrix-Variate [Shv15]. Max [Lyu10]. Max-Type [Lyu10]. Maxima [Khi14b, Kuz11c, Shk11b, ZKK12, ZKK13]. Maximal [AGGR15, Leb10, Mak17, Mum11, Mum12, Yon12]. Maximization [Gus11, Iva15, Mor12, Vos17]. Maximum [Afa14, Bor15, CLS15, Koz14, MS19, Mur11, NN11, Roz10, Sav15, SW11, WN11]. Maxwell [Moh14]. McKean [But14]. Mean [GH11, Pin19a]. Means [GH14]. Measurability [Kol12]. Measurable [Bak15]. Measure [BO13, NBS13, Rok10, Yak19]. Measures [Bak15, CJJ11, Fat14, Fel12, Kaj17, Kon10, Kon11b, KPZ18, MS13, PST17, Rad10, Rad15, Rad18, Sha12c, Su15]. Mechanics [Shi18a]. Medium [Lag13b]. Members [Yak10]. Memoriam [Kha14b, Kha14c, Kha15]. Memory [Chi19, Hol16, JS15, Shi19b, VZ16, Yas16, Ibr12]. Merging [Faz12, Faz13]. Meta [BNP12]. Meta-Times [BNP12]. Method [Fat14, ISF14, KP19, KL14, MS19, Pit18].

Methods [Boy12, Kud19, Ali18a, Ali18c, SPTS17, SP19]. Microscopic [LMM13]. Middle [Yak10]. Miller [AL17a]. Minimal [Dar14, Gri10, Pin18, Roz18b, Zai19]. Minimax [BK13b, Shi17a]. Minimaxity [Pol18]. Minimum [Leb19, RMC14]. Miscellanea [GvZ13, PF12]. Mises [Vol11]. Mixed [ADES19, CK19, She18a]. Mixing [BK12, BK13a, KS18, KV13, Mag10, Sid10, WH19]. Mixtures [Il’10, YB11]. Model [AS16, BM16a, Iva15, Kas12, KL11, Lag13b, Sak11, WH19, Zhu14, BBP18]. Modeling [CPC13, Maz18]. Models [AR13, CV10, GM15, HM12, HM13, IS12, RRV16, Rok12, SU11, SCN12, SCN13, Vos17, Yar11]. Moderate [Bor19, CL10]. Moderately [BM14b]. Modern [KS16, KS17]. Modigliani [AL17a]. Modulated [PS10c]. Moment [AP18, Aly17, BX16, GNT18, KS11, KD18, LG11, Lyk18, Roz18b, She13, She18b, SL13, WD16, WHV18]. Moment-Type [KS11, She13]. Moments [KP19, KP12, KP13, Lyk18, Nag12, NKL14, Pit18]. Monotone [PG14]. Monte [Kud19, Moh14]. Morton [Rok10]. Moscow [Shi18a]. Most [Nov11, NN12]. Motion [AL19, GP17, GU16, Ma13, Mur11, NKL14, Shi17a, Sin12, Spi12, Vak12, ZM13, ZS14]. Moving [DSW19, HM12, HM13, Sid10]. Multichannel [AT14, Bel12, Kry19]. Multicomponent [MR15]. Multidimensional [NR16, Rag10, Zor15]. Multiparameter [Ker16]. Multipath [Asg14]. Multiple [BB19, Bys15, FR13, Yak16, Yas11]. Multiplicative [Su15]. Multiscale [SS17]. Multiserver [Gri19]. Multitype [Dya12, GIKP15, Vat11c, VD18b, VD18a]. Multivariate [Kas12, Leb13, Ruk10, Sak19, Sen10b, Tyu11b, Yam12].

N [Kha10, Kha16b]. Naive [Rok19].

National [CDD15]. Natural [Kru10]. Near [Pol18]. Near-Minimaxity [Pol18]. Nearly [GIKP15]. Necessary [Bah10, Zai15]. Negative [Bor15, Sin12, VUC13, WN11, Yak11]. Negatively [Nez11, Wu15]. Neighborhood [Bul16]. Networks [KME18]. News [Bor12, Bor13, Bul18, Dor10, Kha12b, Kha13b, Kha17a, Kha18b, KMS14, Man15, Mol12, Nik13, PS14, PS10a, PS10b, She14a, Shi11, Shi15b, Shi15a, Shi18a, Su13, Vat11b]. Nikolai [Kha15]. No [BRU10, Sav19]. No-Loss [BRU10]. NOD [WHV18]. Noise [AB16b, BK12, BK13a, Hol18, KPP14]. Noncentral [EST11]. Noncoincidence [PST17]. Noncollinearity [Pin18]. Nonhomogeneous [CPC13]. Nonidentically [Bor10]. Nonimprovable [Sen13]. Nonlattice [Roz19a]. Nonlinear [BKS18, But14, LB18, Rag10, RA13]. Nonnegative [Gus18, Roz14]. Nonorthogonal [Yas11]. Nonparametric [Ale10, NBS13, WH19]. Nonrandomized [FP10, Zai15]. Nonruin [BK16]. Nonstationary [Sta15]. Nonsymmetric [Bor14]. Nonuniform [NS13, Sen11, Sen15, Syu10b]. Norm [NP16, Pus11, Roz19b]. Normal [Bor10, LG11, NC12, PGU18, She13, She14b, Su15]. Normality [Sha15]. Normalization [PG14]. Note [HZ10, Mag10, PMS19, Prz15, Var15]. Nowak [KPZ18]. Number [BN12, Gri19, IMV15, Moh14, SF11, Yak15, Zhu14]. Numbers [ACS17, Chi16, Gap11, Sen10a, Yar11, Yas10, Zhu14].

O. [Chi19]. Obituaries [Kha15]. Observation [Kas12]. Observations [ADES19, NN12, Vol11]. Offord [GEZ18]. Olkin [BNB18]. One [Ark15, BN19, BE14, Fel12, Gam18, GP17, Hol19, ISF14, Irk11, Kon11b, KS15, Kru14, Lin16, Lin18, NC12, NN12, Sen12, Sta15, Zhu11]. One-Dimensional [Ark15, BE14, Hol19, ISF14].

One-Parameter [KS15]. One-Sided [NN12]. One-Step [Lin16, Lin18]. Only [HTV12]. Operator [ISF19a]. Operators [GS12, Hol17, ISF19b, Jaj10, Pla17]. Optimal [Ark15, AB16b, BBD11, Bah10, BZ17, BRU10, BP16, BM16a, CM17, Dem14, FR13, HY16, Pin19b, PS10c, RRV16, Sak11, Sin12, ZS14]. Optimality [Bah10, Shi17a, TPP12]. Option [CV10, Vos17]. Optional [PN15]. Options [NAKL17]. Order [Ale10, Bol13, BV16, Bor15, GH14, KS15, PG14, Rod19, Shi19a, Yak18, TPP12]. Ordered [Yak10]. Ordering [BNB18]. Origin [HTV12]. Ornstein [Sza13, ADES19, BCS11, BD13, CK19, LS19, Pal19, Sza12, Vak12, dRdC17]. Orthant [Yak19]. Our [Shi17b]. Outer [Var19]. Overtake [HM12, HM13].

Pairwise [Wu15]. Parabolic

[AL19, Bel17, Gla16, Rai13, Sha12c].

Parameter

[ADES19, GM15, Kol12, KS15, Shu14, VP10]. Parameter-Measurability [Kol12]. Parameters [BP16, Bor18b, Sta15]. Partial [BB19, Roz10, Yon12]. Partially [KMM15]. Particle [Che15, Kar17]. Particles [HM12, HM13, Kon11a, Leb13, VL13, Yar11, Zhu14]. Partitions [Su15]. Pass [Kon10]. Passage [Bor17b, DSW19, Kaj17]. Path [Che15, Zan13]. Paths [BN12]. Pathwise [Sau19]. Pattern [Ruk10]. Payoff [FR13]. Pecherskii [Kuz11a]. Penalization [NN11]. Penetration [Lag13b]. Performance [YYZ13]. Periodic [PR19]. Periodically [MR15]. Permutation [Yak18]. Permutations [Tim17, Yak10]. Persistence [AM19b]. Perspective [CKT16]. Perspectives [VK16]. Perturbed [AB16b, Koz17]. Phase [GM15]. Physical [BK12, BK13a]. Pitman [NK13, NKL14]. Plachky [Com18]. Planar [Vak12]. Planck [BKS18, Sha12b]. Point [Asg14, Dar12, Dar14, Pol18, SS17]. Points [Pin18]. Pointwise [Jaj10]. Poisson [Bak10, BBP18, BK13b, HZ11, She14b, She18a, She18b, WN11]. Poisson-Binomial [HZ11, She18b]. Pollak [Pol18]. Polling [Vat11c]. Pòlya [Ruk10]. Polyharmonic [Kel10]. Polynomial [Osm10, SF11]. Polynomials [GZ12, GZ13, Sen10b, Uly16]. Populations [LV19]. Porous [Lag13b]. Portfolios [BNB18, CM17]. Portmanteau [GP10]. Positive [PMS19, Roz16, Roz18a]. Possible [PF12]. Posteriori [Kry16]. Potential [Bak15]. Power [Iva15, Nag12, Roz16, Roz18b, Tim18]. Power-Series [Tim18]. Powerful [Nov11, NN12]. Powers [PGU18]. Precision [Sen16]. Predictable [TE16]. Premium [Irk11]. Price [Kha13a, Rok12]. Prices [CV10]. Pricing [CKT16, NAKL17]. Prime [Tim17]. Principle [Afa10, Afa11b, Bor14, Irk11, IS12, YYZ13]. Principles [BM12b, BM13a, BM13b, BM14a, BM14b, BM16b, BM16c, Bor19]. Prior [Dar14]. Prize [Yas15]. Probabilistic [Boy12, ISF14, ISF17b, ISF18, Nag15b, Pla17, Rai13]. Probabilities [AB16a, Aib17, AA18, AM19b, BK16, FP10, FKZ14, KP19, Kry16, Mas12, MPR10, Rag10, Roz10, Roz15, Roz16, Roz18b, Roz18a]. Probability [BL12, Fel19, Gre10, Il’10, KPZ18, Lot15, MM13, PST17, Pin19b, Rao18, Sha12b, Shi18a, Sok14, SL13, VD18a, VD18b, Zai15, Zha12, Zhd16, ZS13a, Zor15, Zor18]. Problem [Afa19, AP18, Ali13, Ark15, BBD11, BZ17, Dar12, GEZ18, Gus11, ISF14, JKV14, Lis19, LG11, Lyk18, Mor12, Pal18, Pla17, Rai13, Sha12b, SL13, ZM13]. Problems [Asg14, Bah10, BCG13, ISF15, ISF17b, ISF18, Iva15, KS16, KS17, Kaj17, LB18, Shi17a, ZS13a, ZS14]. Procedure [Pol18]. Procedures [Kar13, Kar14, TPP12]. Process [Afa10, Afa11b, ADES19, Ali13, AM19b, AGGR15, BCS11, BN12, BK13b, CL10, CK19, FV10, GD15, Har16, Kas16, Kel10, Khi10, Koz14, LS19, Mal18b, Mas12, Pal19, Sta15, VW10, VL13, Yas10, ZX19, ZK18, dRdC17]. Processes [AKG16, Ark15, AM19b, BRKF11, BB19, BD13, BM14b, BM16b, BM16c, Bor18a, Bor19, BM16d, Bul15, BS18, Bys15, CO11, Cor11, DP13, Dya12, EK19, EST11, Faz12, Faz13, GGW14, Gla16, GU16, GIKP15, HKP16, ISF17a, ISF19b, Kaj17, Khi14a, KP19, Kor17, Koz10, KPH14, Kru10, Kru12, Kru13, Kud19, LaG13a, Leb10, Leb13, LV19, Mak17, Mam12, Mam13, MR15, Nag15b, NP16, NN11, Noa10, Nov11, PP10, Pus11, SV18, Sau19, Sav19, Shi19a, SK11, SP18, Sza12, Sza13, Top10, Vak12, Vat11c, Vat15, Vat16, VD18a, VD18b, Vos17, Zan13, Zhd16, ZP18, vDZP10]. Product [FV10, Vat11c, Zhd16]. Products [GS12]. Progeny [Koz10]. Prokhorov [GvZ13, Hol16, Kha14c, Pro18, Vat11a, VK16]. Proof [Kru12, Kru13, Kuz11a, Nag10, ZS13b]. Properties [ACS17, BKNP15, But14, CV10, ISF17a, Kor17, Leb13, Lin18, MPR10, Sha12c, SCN12, SCN13, Spi12, Uly16, Yon12]. Property [Cor11, CJJ11, Gam18, Mur11, TE16, Var15]. Proximity [NC12]. Pseudogolden [BO13]. Pseudoweak [GP10]. Pulse [KPP14].

Quadratic [BN19, BKNP15, Pus11, Wu15]. Qualitative [Sha12c]. Quantifying [Pin18]. Quantity [CS17]. Quantum [Gre10, Hol18, Hol19, Kuz11b]. Quasi [PMS19, Sen11, Syu10b]. Quasi-Nonuniform [Sen11, Syu10b]. Quasi-stationary [PMS19]. Quasilinear [Bel17]. Quaternion [VC10]. Queue [AB16a, AA18]. Queueing [AR13, AT14, AT19, Bel12, Gri19, Koz14, Kry19, SZK13].

R. [Shi19b]. Radial [dRdC17]. Radius [Bul16, Tim18]. Ramachandran [Ibr13]. Random [Afa10, Afa13, Afa14, Afa17, Afa19, AM19a, Aly17, Bak10, Bar13, BBE17, BS14, BP15, BKNP15, BN12, Bor14, Bor10, BM12a, BM12b, BM13a, BM13b, BM14a, BM14b, Bor15, Bor17b, Bor17a, Bul11, Bul16, Bur16, CO11, Che15, Che11, CL10, Dar14, DP13, DP19, Dem14, Dor10, Dya12, EZ13, Fel17, Gam15, Gol13, GZ12, GZ13, GNT15, GM15, GM19, GP10, Gri19, HTV12, ISF15, ISF19a, JSS16, KS18, Kha18a, KY19, Kor14, KD18, Koz10, Koz14, Kru14, Kuz11c, KT14, Lau11, Leb13, Leb19, LMM13, Lyk18, Lyu10, Ma17, Mal18a, Mal17, Mam12, Mam13, MS13, Moh14, Naz10, Nez11, ORS14, PP17, Pin18, Pin19b, PR19, PS10c, Rag10, Roz14, Roz15, Roz16, Roz18b, Roz18a, Roz19b, Roz19a, RMC14]. Random [Sav15, Sem19, Ser10, Ser11, Sha15, Sha16, SF11, She13, She18b, Shk11b, SW11, Shu17, Son16, Su15, Tim17, Tim19, Tyu10, Uly16, VC10, VW10, Vat11c, VL13, VD18a, VD18b, WD16, WH14, WHV18, WH19, Wu15, Yak10, Yak15, Yak18, Yar11, Yar18, Yas11, YB11, Yon12, Zhu11, ZKK12, ZKK13]. Randomization [Zai15]. Randomized [Pol18]. Ranking [Kar14]. Rarefied [Kor17]. Rate [BK12, BK13a, Gap11, GT10, Kar17, KS11, KP12, KP13, NS13, Sak11, SZK13, She18a, She18b, Tyu11a, Yas10, vDZP10]. Rates [PL10]. Ratio [BO13, Shu11, Shu13]. Reaching [Afa13, Afa17]. Readers [Shi17b]. Real [BOGP14, SF11]. Rearrangements [CLS15]. Recurrence [Rok12]. Recurrent [Shu17]. Recursive [Gol14]. Reduced [LV19, Vat15, Vat16]. Reducibility [Irk11]. Refinement [Bor17a, Kuz11c, Lin16, Tyu12, Tyu13]. Refinements [Sok14]. Reflected [HO16, Shi19a]. Reflecting [AKG16, ISF19b]. Regenerative [AT14, AB16a, AT19, AA18, Kry19].

Region [KP19]. Regression

[KPP14, KS15, Lin18, LB18, Mum11,

Mum12, NBS13, WH19]. Regular [ORS14]. Regularity [Kol13, Mal18b, Sha12c]. Regularly [Yak19]. Related [CO11, KT14]. Relation [Kon11b]. Relations [Rok12]. Relationships [Ma13]. Relative [Kha18a]. Relaxed [KD18]. Remainder [Tyu12, Tyu13]. Remarks [BK17]. Remote [Bor17b]. Renewal [BM16b, BM16c, Bor18a, Bor19, BS18, Kor17, Nag12, Nag15a, VT14]. Renko [Spi12]. Replacement [Tol17]. Replacements [ZK18]. Representation [GP17, JGM16, Pla17, Rai13, TE16]. Representations [ISF17b, ISF18, Kel10, Lyu10, PP10]. Rescaled [FV10]. Residuals [KS15, Syu10b]. Respect [Bys15, GP10, Kol12, YYZ13]. Result [RMC14]. Results [BNB18]. Reviews [Kho13]. Revisited [KPZ18, NK13]. Riemann [Boy12, Pla17]. Rigidity [GNT18]. Risk [Kon10, Kon11b, Zai19, ZX19]. Roberts [Pol18, TPP12]. Robust [Bol19, Gus11, Mor12]. Robustness [Bol13]. Rogozin [Kuz11a]. Root [Bol19]. Roots [GZ12, GZ13]. Rosenblatt [AM19b]. Rosenthal [Pin15]. Rosenthal-Type [Pin15]. Rotations [Mal17]. Rowwise [Wu15, Yon12]. Ruin [KL11, Lot15]. Rule [GL13]. Russian [CDD15, SU11, Yas15].

Sample [Bao15, GT10, Kar13, Kar14, ORS14, Rad10, Yas18, Zai15, Zan13]. Samples [BN19, Tol17]. Sampling [Moh14, PN15]. Satisfying [PST17]. Scaling [Che15, Cor11]. Scheme [BBD11, Bor10, MZ18]. Schoenberg [Ma13]. Science [CDD15]. Scientific [Bor12, Bor13, Bul18, Dor10, Kha12b, Kha13b, Kha17a, Kha18b, KMS14, Man15, Mol12, Nik13, PS14, PS10a, PS10b, She14a, Shi11, Shi15b, Shi15a, Shi18a, Su13, Vat11b].

SDE [KV13]. SDEs [HO16]. Second [Bor15, GH14, KP12, KP13]. Segment [ZS14]. Selection [GL13, Kar13]. Self [Ali13, AS16, EST11]. Self-Exciting [Ali13]. Self-Similar [AS16, EST11]. Selfdecomposable [CJJ11]. Semenovich [Kha14a]. Semi [Har16, Kas16]. Semi-Markov [Har16, Kas16]. Semicircle [GNT18]. Semimartingale [Khi14a, Shi19a]. Semimartingales [Khi10]. Seminar [Hol16, Shi18a]. Semistable [Faz12, Faz13]. Semistationary [Sau19]. Sense [Fel12]. Sequence [Dar12, Dar14, Khi14b]. Sequences [AL17b, BS11, Gap11, Gol14, Gri10, Nez11, PS10c, Sid10, Tyu10, ZKK12, ZKK13, ZN11]. Sequential [BP16, Kar13, Kar14, Lis19, LS19, Nov11, NN12, YMW15]. Series [Maz18, Nag10, Rad18, Roz13, Roz14, Tim18, Yak16, Yas18]. Server [Aib17, AA18]. Servers [Gri19]. Service [Gri19]. Set [Kas16, Naz10, Ser11, Tim19]. Sets [GP10, Sha16]. Sevast’yanov [VZ16, Kha14b]. Sh [Shi19b]. Shannon [Kha16a]. Shape [HKP16, KPH14]. Sharp [BCS11, BC10]. Shepp [Shk11a]. Shiryaev [JKV14, Kha10, Kha16b, Pol18, PMS19, TPP12]. Short [Bor19, HZ10]. Siblings [DP18a]. Sided [NN12]. Sieve [IMV15]. Sign [Bol13, Bol19]. Signal [AB16b]. Signs [CLS15]. Similar [AS16, EST11]. Similarity [BNHS17]. Simple [DP13, Kru12, Kru13, NN12, Sav19]. Simplest [BN12, Bor14]. Simultaneous [Gri19]. Single [AA18, Moh14]. Single-Server [AA18]. Singular [MPR10]. Singularities [Asg14]. Size [Kar13, Kar14, Zai15]. Sizes [Tim19]. Skew [Sak19]. Skitovich [Ibr13]. Skorokhod [Ibr12, FV10, JKV14, LR10]. Slightly [GH14]. Slowly [Lin16]. Small [AL17b, Bur16, LV19, NP16, NP13, Pus11, Roz10, Roz14, Roz16, Roz18b, Roz18a, Roz19b, VP10, ZG17]. Smooth [BS14, Pin15, Roz14]. Sobolev [Kol13, Li15]. Solenoids [Fel10]. Solution [BM16a, BK12, BK13a, Pal18, Pla17, Rai13, Sha12b]. Solutions [BK16, DP18b, ISF15, ISF17b, ISF18, LG11, Sha12c, Sha12a]. Solvability [BKS18]. Solving [ISF14]. Some [BK17, Bor14, Fel18, Gap11, Lag13b, NSB10, NP13, Noa10, Roz19b, Spi12, Wu15, Yan12, Yan13]. Sources [KY19, PR19]. Space [Fat14, FV10, Mal17, Mam12, Mam13, Mam17, PST17, Rad15]. Spaces [AS16, Bak15, CO13, Jaj10, Pin15, Rad10, ZS13a]. Spectra [GT10]. Spectral [Ale10, Bak15, BX16, Yar18]. Spectrum [Yas18]. Spent [Lyu12]. Spheres [CO13]. Spine [Mal18a]. Spitzer [Nag10, Roz13]. Square [NBS13]. Squared [dRdC17]. Squares [Zan13]. Stability [AT19, BP15, BCG13, Koz17]. Stable [Cor11, DSW19, EK19, Mal18a, Maz12, Pal18, Shi19a, Syu14, TE16, WD16]. Stable-Like [Shi19a]. Stackelberg [Rok19]. State [AA18, Lyu12, Shi18a]. States [Kas16]. Stationary [AL17b, BS11, BKS18, Gap11, Gol13, Gri10, Hol18, KS18, Leb10, LS19, MR15, Noa10, Zan13, Zhd16, PMS19]. Statistic [Shi17a, WF18]. Statistical [ACS17, CK19, Kry19, Sav17, SU11, Tyu11b]. Statistics [Bao15, BX16, BP15, BL12, BV16, Fel10, JSS16, KMM15, KS15, MF15, PG14, PP17, Rod19, Shk11a, Vol11, Yan12, Yan13, BBP18]. Steady [AA18]. Steinebach [Com18]. Step [Lin16, Lin18]. Stieltjes [GNT18]. Stitched [BM16d]. Stochastic [Ali18a, Ali18c, AKG16, BOGP14, Bel17, BP16, BE14, But14, Bys15, DP18b, Dar12, Dor10, EST11, GS19, Gol14, HY16, Kar17, Kha18a, KL14, Kol12, Kru12, Kru13, LR10, Lyu10, Mam12, Mam13, Mam17, MZ18, Pal18, PP10, Rad10, Rad15, Rad18, Rok19, Sha12a, SU11, SPTS17, SP19]. Stochastically [Kor17]. Stone [Bor17a]. Stopping [Ark15, BBD11, BZ17, BRU10, FR13, PS10c, Sin12, ZS14]. Strategies [Ark15, CS17, IS12, KPZ18, Spi12]. Strictly [Koz10]. Strip [BN12]. Strong [CPC13, Chi16, Gap11, Kol12, Maz18, Sha12a, Shu11, Shu13, Yas11]. Strongly [FP10]. Structure [Kas16, Khi14b, She13, Vat15, Vat16, WH14]. Student [Sav15]. Studies [VK16]. Sub [AS16]. Sub- [AS16]. Subcritical [Koz10]. Subexponential [AA18]. Subexponentiality [Nag11]. Subexponentially [Pal18]. Subgraphs [Bur16]. Submartingale [Gus18]. Submartingales [PN15]. Subordination [BNP12]. Success [Zai15]. Sudden [VW10]. Sufficient [Bah10, Irk11, KMM15, Nag11]. Sum [BB19, Che11, FR13, Fel17, Lyk18, Roz16, Roz18b, Roz18a]. Sum-Payoff [FR13]. Summands [BCG13, CLS15]. Summation [GM15]. Sums [AL17b, Bar13, Bor10, BM12a, Bor17a, EZ13, ISF15, ISF19a, Kha18a, Kor14, KD18, KS15, Kru14, Kuz11c, Pin19b, Roz10, Roz15, Roz19a, She13, She18b, Yas11, Yon12]. Sup [SP18]. Sup-Functionals [SP18]. Supercritical [KY19, Leb13, Yar18]. Superdiffusion [AS16]. Superlarge [Roz15]. Support [GM19]. Supports [Fel12]. Suprema [Sak15]. Sure [Var19, ZN11]. Surfaces [MS13]. Survey [Kov14]. Survival [VD18a, VD18b]. Survive [VD17]. Switching [HY16, Sav19]. Symmetric [EK19, GGW14, SF11, She13]. Symmetrized [Gam18]. Symposium [SU11]. Synchronization [Kar17]. System [Aib17, Bel12, Gri19, Kon11a, Kry19, Rai13]. Systems [AB12, AB13, AR13, AT14, AT19, Bah10, Bel17, Dor10, HY16, Kar17, Kuz11b, LMM13, SZK13, Vat11c].

Tail [AA18, Leb10, Pin19b, SCN12, SCN13, VT14]. Tailed [Nov14, SF11, GZ12, GZ13]. Tails [BC10, BD13, Bys15, Maz18, Rod18, Rod19, Roz15]. Talks [Ali18a, Ali18c, SP19].

Tanaka [EK19]. Tau [Leb19]. Tauberian [Yak16]. Technology [CDD15]. Telegraph [DP18b]. Tempered [BRKF11]. Tensor [Sak10, Sak11]. Terminal [Gus18]. Test [BN19, Bol19, Rod18, Zai15]. Testing [Bol13, Lis19, LS19, Sak19, Zai15, ZM13]. Tests [Bol13, NSB10, Nov11, NN12, SP18]. Their [AR13, BNB18, CM17, DP18b, HM12, HM13, Kor17, NN11, NR16, VK16, WHV18, Yam12]. Them [Khi14b]. Theorem [AL17a, Afa14, AKG16, Bao15, Bor17a, Bul17, CKT16, EST11, Fel18, Fel19, Gam15, GP10, Ibr13, KS18, Ker16, KY19, KS12, KS11, KPZ18, Kuz11c, Mag10, Nag12, PN15, Rok10, Sav15, Sen15, Sha16, Tyu11a, Tyu12, Tyu13, Var19, VT14, Yak10, Yak16, Yak19, ZK18, BS11, Com18, KP12, KP13, MS19, MS13, NS13, Sen11, Sen12, Sen13, SS13]. Theorems [AB12, AB13, Bor10, Bor18a, BS18, DP13, DP19, GNT15, GM15, GM19, Gri10, ISF15, ISF17a, ISF17b, ISF18, Jaj10, KH14, KH15, Nag15a, Nez11, RA13, Shk11b, Shu11, Shu13, Shu14, Tim18, Vat16, Vol11, Wu15, Yan12, Yan13, ZN11]. Theory [Boy12, Cor11, Kry16, MM13, Tyu11b]. Third [BV16, TPP12]. Third-order [TPP12]. Three [Mal17]. Three-Dimensional [Mal17]. Threshold [Ark15]. Tilted [Pin19a]. Time [Afa13, Afa17, AL19, Bor17b, DP13, Deh13, Gre10, Gri19, Lyu12, Maz18, Nov11, Pal19, Pol18, PR11, RA13, Rok12, Son16, Top10, VD17, Yas18, ZKKS17, ZP18]. Time-Varying [Pal19]. Times [BNP12, DSW19, GvZ13, Vak12]. Timing [CS17]. Toeplitz [GS12]. Topological [AS16]. Total [Sav17]. Trace [GS12]. Trading [Spi12]. Traffic [AR13, Kry19, SW11]. Trajectories [BM12a, BM12b, BM13a, BM13b, BM14a, BM14b, BM16b, BM16c, Bor19, HKP16, KPH14]. Trajectory [Kas16]. Transfer [Ker16]. Transform [GNT18].

Transformations [Bor14, Naz10]. Transforms [Zha12]. Transient [Afa17]. Transition [FP10, Hol17, MPR10]. Transitions [GM15]. Transport [Kol13]. Transportation [Kol13]. Triangular [Bor10]. Tributes [Ali18b]. Trimmed [GH11, GH14]. True [KL14, PR11]. Truncated [GS12, Pin19a]. Truncated-Tilted [Pin19a]. Truncation [ZKKS17]. Truncations [SZK13]. Turbulence [BNHS17]. Two [Afa19, BNB18, BKNP15, Bul11, GNT15, GM15, Kol12, LS19, Lyk18, Sem19, Shu14, Sta15, VL13, Zhd16]. Two-Boundary [Afa19]. Two-Colorings [Sem19]. Two-Dimensional [Bul11, Sta15]. Two-Parameter [Kol12]. Type [BCG13, BM12a, BD13, CL10, DP19, KPP14, KS11, Lyu10, NSB10, NAKL17, PL10, Pin15, Rod18, She13, WN11, GGW14]. Types [VL13].

Uhlenbeck [ADES19, BCS11, BD13, CK19, LS19, Pal19, Sza12, Sza13, Vak12, dRdC17]. Unbounded [MM13]. Unboundedness [Kru12, Kru13]. Uncertain [Dar12]. Unified [LR10, SW11, YYZ13]. Uniform [Gol13, HZ11, Shu11, Shu13, Tyu10]. Uniformly [Zai19]. Uninformed [Rok19]. Uniqueness [Sha12a, Sha12b, Kra16]. Unit [Bol19]. Units [Kas12]. Universal [ZS13b]. University [Shi18a]. Unknown [BP16]. Unreliable [Aib17]. Until [Lyu12]. Upper [Kha13a, KS10, Leb19, Pal19, Pin19b]. Using [LR10, Moh14, Rod19]. Utilities [CM17]. Utility [Gus11, Iva15, Mor12, Vos17].

V [Bin14, Chi19, Hol16, Kha14c, PSM14, Vat11a, VK16]. V. [Kha13c]. Vakhania [Kha15]. Validity [Per12]. Value [Gus11, ISF14, ISF15, ISF17b, ISF18, Mor12, Rai13]. Values [Gus18, VP10]. Variable [Aly17, Lyk18, Shi19a, ZG17]. Variables [Bar13, BP15, BKNP15, Bor10, Che11, Dem14, EZ13, Fel17, GM15, ISF15, ISF19a, Kha18a, Kru14, Lau11, Leb19, Lyk18, Nez11, Pin19b, PGU18, Roz14, Roz15, Roz16, Roz18b, Roz18a, Sav15, She13, VC10, WHV18, WH19, Wu15, Yas11, Yon12, Zor15, ZKK12, ZKK13]. Variance [Bor15, KL11, NK13]. Variate [Shv15]. Variation [EST11, Sav17]. Variations [KS12]. Varying [KY19, Pal19, Yak19]. Vasilyevich [GvZ13]. Vatutin [Kha13c]. Vector [AM19a, KP19, Ma17, Shv15]. Vectors [BKNP15, BM12a, Bor17a, Gam15, Roz19a]. Velocities [GS19]. Vershikian [Lau11]. Version [Ker16]. Versions [WN11]. Very [Zor15]. via [NAKL17, Vak12]. Viscosity [BK16]. Visits [PP17]. Viskov [Chi19]. Vladimirovich [Ibr12, Kov14]. Vlasov [But14]. Volatility [BP16, ZG17]. Volume [HZ10]. vs [NN12].

Walk [Afa13, Afa14, Afa17, Afa19, BM12b, BM13a, BM13b, BM14a, Bor15, Bor17b, Bul11, Che15, HTV12, Koz14, Lyu10, Mal18a, PP17, Rag10, Shk11b, WD16, WH14, Yar18]. Walks [BN12, Bor14, BM12a, BM14b, CO11, DSW19, KY19, PR19, SW11, Shu17, Yar11, YB11]. Wang [Irk11]. Warming [PF12]. Watson [Afa11b, Dya12, GIKP15, Nag15b, Top10]. Wavelet [RMC14]. Weak [AKG16, CPC13, FV10, Gri10, HS16, IMV15, JGM16, MZ18, Var15, Var19, Yas10, Zhu11]. Weakened [PST17]. Weakly [FKZ14, Sha15, Maz12]. Weibull [Rod18]. Weibull-Type [Rod18]. Weighted [EZ13, Lin18, Pus11, Roz16, Roz18b, Roz18a, Sak15, SP18]. Weights [Roz14, Yak18]. Which [Syu14]. White [BK12, BK13a]. Wiener [Sza12, BN12, Kud19, Kuz11a, Sza13]. Wigner [BX16]. Willinger [Rok10].

Wilson [BM16d]. Winding [Vak12]. Wins [Yas15]. Withdrawal [ZG17]. within [WH14]. without [HM12, HM13, Tol17]. Wolfowitz [NP16]. Work [Bin14]. Works [Kov14]. Worldwide [Bin14].

Yor [JS15]. Yu. [Hol16, Kha14c, Vat11a, VK16]. Yuri [GvZ13, Pro18]. Yushkevich [Mol12].

Zero [Fel12, Kuz11c, Roz16, Roz18b, Roz18a, SK11, Zhu11]. Zero-One [Fel12, Zhu11]. Zeros [SF11]. Zeta [Boy12]. Zeta-Function [Boy12].

References

Aibatov:2018:SAS

[AA18] S. Z. Aibatov and L. G. Afanasyeva.

Subexponential asymptotics for steady state tail probabilities in a single-server queue with regenerative input flow. Theory of Probability and its Applications, 62(3):339–355, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanaseva:2012:LTS

[AB12] L. G. Afanas’eva and T. N. Belorusov. Limit theorems for systems with impatient customers under high load conditions. Theory of Probability and its Applications, 56(4):674–682, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanaseva:2013:LTS

[AB13] L. G. Afanas’eva and T. N. Belorusov. Limit theorems for systems with impatient customers under high load conditions. Theory of Probability and its Applications, 56(4):674–682, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanasyeva:2016:PLD

[AB16a] L. G. Afanasyeva and E. E. Bashtova. Probabilities of large deviations for a queue with regenerative input flow. Theory of Probability and its Applications, 60 (1):120–126, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Artemov:2016:OES

[AB16b] A. V. Artemov and E. V. Burnaev. Optimal estimation of a signal perturbed by a fractional Brownian noise. Theory of Probability and its Applications, 60 (1):126–134, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Avdeeva:2017:ESP

[ACS17] M. Avdeeva, F. Cellarosi, and Ya. G. Sinai. Ergodic and statistical properties of B-free numbers. Theory of Probability and its Applications, 61(4):569–589, ???? 2017. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Alazemi:2019:BEB

[ADES19] F. Alazemi, S. Douissi, and Kh. Es-Sebaiy. Berry–Esseen bounds and ASCLTs for drift parameter estimator of mixed fractional Ornstein–Uhlenbeck process with discrete observations. Theory of Probability and its Applications, 64 (3):401–420, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanasyev:2010:IPC

[Afa10] V. I. Afanasyev. Invariance principle for the critical branching process in a random environment attaining a high level. Theory of Probability and its Applications, 54(1):1–13, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanasyev:2011:BHJ

[Afa11a] V. I. Afanasyev. Brownian high jump. Theory of Probability and its Applications, 55(2):183–197, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i2/p183_s1.

Afanasyev:2011:IPC

[Afa11b] V. I. Afanasyev. Invariance principle for a critical Galton– Watson process attaining a high level. Theory of Probability and its Applications, 55(4):559–574, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i4/p559_s1.

Afanasyev:2013:ATR

[Afa13] V. I. Afanasyev. About time of reaching a high level by a random walk in a random environment. Theory of Probability and its Applications, 57(4):

547–567, ????  2013.  CO-

DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanasyev:2014:CLT

[Afa14] V. I. Afanasyev. Conditional limit theorem for the maximum of a random walk in random environment. Theory of Probability and its Applications, 58 (4):525–545, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanasyev:2017:TRH

[Afa17] V. I. Afanasyev. On the time of reaching a high level by a transient random walk in a random environment. Theory of Probability and its Applications, 61 (2):178–207, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanasyev:2019:TBP

[Afa19] V. I. Afanasyev. Two-boundary problem for a random walk in a random environment. Theory of Probability and its Applications, 63 (3):339–350, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aydogmus:2015:CMB

[AGGR15] O. Aydogmus, A. P. Ghosh, S. Ghosh, and A. Roitershtein. Colored maximal branching process. Theory of Probability and its Applications, 59(4):663–672, ???? 2015. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aibatov:2017:LDP

[Aib17] S. Aibatov. Large deviation probabilities for the system M/G/1/∞ with an unreliable server. Theory of Probability and its Applications, 61(2):321–326, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aliyev:2016:WCT

[AKG16] R. Aliyev, T. Khaniyev, and B. Gever. Weak convergence theorem for ergodic distribution of stochastic processes with discrete interference of chance and generalized reflecting barrier. Theory of Probability and its Applications, 60(3):502–513, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aboura:2017:NDM

[AL17a] S. Aboura and E. Lepinette. New developments on the Modigliani– Miller theorem. Theory of Probability and its Applications, 61 (1):3–14, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aurzada:2017:SDS

[AL17b] F. Aurzada and M. A. Lifshits.

Small deviations of sums of correlated stationary Gaussian sequences. Theory of Probability and its Applications, 61(4):540–568, ???? 2017. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aurzada:2019:FET

[AL19] F. Aurzada and M. A. Lifshits.

The first exit time of fractional Brownian motion from a parabolic domain. Theory of Probability and its Applications, 64(3):490–497, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Alekseev:2010:NEH

[Ale10] V. G. Alekseev. On nonparametric estimates of higher order spectral densities. Theory of Probability and its Applications, 54(3):499–504, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i3/p499_s1.

Aliev:2013:DPS

[Ali13] A. Aliev. Disorder problem for self-exciting process. Theory of Probability and its Applications, 57 (3):512–520, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Alimov:2018:ATG

[Ali18a] A. Alimov. Abstracts of talks given at the 2nd International Conference on Stochastic Methods. Theory of Probability and its Applications, 62(4):640–674, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). See erratum [Ali18c].

Alimov:2018:BT

[Ali18b] A. R. Alimov. Birthday tributes. Theory of Probability and its Applications, 62(1):173, ???? 2018.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Alimov:2018:EAT

[Ali18c] A. R. Alimov. Erratum: Abstracts of Talks Given at the 2nd International Conference on Stochastic Methods. Theory of Probability and its Applications, 63(2):337, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). See [Ali18a].

Aly:2017:MEA

[Aly17] S. M. Aly. From moment explosion to the asymptotic behavior of the cumulative distribution for a random variable. Theory of Probability and its Applications, 61(3):357–374, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Alsultan:2019:DVR

[AM19a] R. Alsultan and C. Ma. Kdifferenced vector random fields. Theory of Probability and its Applications, 63(3):393–407, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aurzada:2019:PPD

[AM19b] F. Aurzada and C. Mönch. Persistence probabilities and a decorrelation inequality for the Rosenblatt process and Hermite processes. Theory of Probability and its Applications, 63(4):664–670, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afendras:2018:FMD

[AP18] G. Afendras and N. Papadatos. A factorial moment distance and an application to the matching problem. Theory of Probability and its Applications, 62(3): 496–504, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanasyeva:2013:QST

[AR13] L. G. Afanasyeva and I. V. Rudenko. GI|G|∞ queueing systems and their applications to the analysis of traffic models. Theory of Probability and its Applications, 57(3):375–395, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arkin:2015:TSO

[Ark15] V. I. Arkin. Threshold strategies in optimal stopping problem for one-dimensional diffusion processes. Theory of Probability and its Applications, 59(2): 311–319, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arkashov:2016:MSS

[AS16] N. S. Arkashov and V. A. Seleznev. On a model of sub- and superdiffusion on self-similar topological spaces. Theory of Probability and its Applications, 60 (2):173–186, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Asgharian:2014:SIM

[Asg14] M. Asgharian. On the singularities of the information matrix and multipath change-point problems. Theory of Probability and its Applications, 58(4): 546–561, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanasyeva:2014:MQS

[AT14] L. G. Afanasyeva and A. V.

Tkachenko. Multichannel queueing systems with regenerative input flow. Theory of Probability and its Applications, 58(2): 174–192, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanasyeva:2019:SCQ

[AT19] L. G. Afanasyeva and A. V.

Tkachenko. Stability conditions for queueing systems with regenerative flow of interruptions. Theory of Probability and its Applications, 63(4):507–531, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bahlali:2010:NSC

[Bah10] S. Bahlali. Necessary and sufficient conditions of optimality for optimal control problems of forward and backward systems. Theory of Probability and its Applications, 54(4):553–570, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bakhtin:2010:PLA

[Bak10] Yu. Bakhtin. Poisson limit for associated random fields. Theory of Probability and its Applications, 54(4):678–681, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bakhtin:2015:SPK

[Bak15] V. I. Bakhtin. Spectral potential, Kullback action, and large deviations of empirical measures on measurable spaces. Theory of Probability and its Applications, 59 (4):535–544, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bao:2015:AEC

[Bao15] Z. Bao. On asymptotic expansion and central limit theorem of linear eigenvalue statistics for sample covariance matrices when N/M → 0. Theory of Probability and its Applications, 59 (2):185–207, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Barseghyan:2013:DSI

[Bar13] A. G. Barseghyan. On distribution of sums of identically distributed random variables. Theory of Probability and its Applications, 57(1):132–135, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borisov:2019:EID

[BB19] I. S. Borisov and A. A. Bystrov. Exponential inequalities for the distributions of canonical multiple partial sum processes. Theory of Probability and its Applications, 64(2):171–185, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Babilua:2011:OSP

[BBD11] P. Babilua, I. Bokuchava, and B. Dochviri. The optimal stopping problem for the Kalman–Bucy scheme. Theory of Probability and its Applications, 55(1):110–119, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i1/p110_s1.

Bedini:2017:BBR

[BBE17] M. L. Bedini, R. Buckdahn, and H.-J. Engelbert. Brownian bridges on random intervals. Theory of Probability and its Applications, 61(1):15–39, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bendikov:2018:PSE

[BBP18] A. Bendikov, A. Braverman, and J. Pike. Poisson statistics of eigenvalues in the hierarchical Dyson model. Theory of Probability and its Applications, 63(1):94–116, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bertail:2010:SBT

[BC10] P. Bertail and S. Clémençon.

Sharp bounds for the tails of functionals of Markov chains. Theory of Probability and its Applications, 54(3):505–515, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i3/p505_s1.

Bobkov:2013:SPC

[BCG13] S. G. Bobkov, G. P. Chistyakov, and F. Götze. Stability problems in Cramér-type characterization in case of I.I.D. summands. Theory of Probability and its Applications, 57(4):568–588, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bercu:2011:SLD

[BCS11] B. Bercu, L. Coutin, and N. Savy. Sharp large deviations for the fractional Ornstein–Uhlenbeck process. Theory of Probability and its Applications, 55(4):575–610, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v55/i4/p575_s1.

Borovkov:2013:OUT

[BD13] K. Borovkov and G. Decrouez. Ornstein–Uhlenbeck type processes with heavy distribution tails. Theory of Probability and its Applications, 57(3):396–418, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Blei:2014:ODS

[BE14] S. Blei and H.-J. Engelbert. Onedimensional stochastic differential equations with generalized drift. Theory of Probability and its Applications, 58(3):345–357, ???? 2014. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belorusov:2012:EMQ

[Bel12]  T. N. Belorusov.  Ergodic-

ity of a multichannel queueing system with balking. Theory of Probability and its Applications, 56(1):120–126, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v56/i1/p120_s1.

Belopolskaya:2017:SIQ

[Bel17] Ya. I. Belopolskaya. Stochastic interpretation of quasilinear parabolic systems with cross diffusion. Theory of Probability and its Applications, 61(2):208–234, ???? 2017. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bingham:2014:WIW

[Bin14] N. H. Bingham. The worldwide influence of the work of B. V. Gnedenko. Theory of Probability and its Applications, 58(1):17–24, ???? 2014. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bondarev:2012:RCS

[BK12] B. V. Bondarev and S. M. Kozyr.

On the rate of convergence of solution of differential equation disturbed by “Physical” white noise to a solution of appropriate diffusion equation. exponential mixing. Theory of Probability and its Applications, 56(4):562–578, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bondarev:2013:RCS

[BK13a] B. V. Bondarev and S. M. Kozyr. On the rate of convergence of solution of differential equation disturbed by “Physical” white noise to a solution of appropriate diffusion equation. Exponential mixing. Theory of Probability and its Applications, 56(4):562–578, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Burnashev:2013:MDP

[BK13b] M. V. Burnashev and Y. A. Kutoyants. On minimax detection of Poisson process intensity. Theory of Probability and its Applications, 57(2):183–195, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belkina:2016:VSI

[BK16] T. Belkina and Yu. Kabanov. Viscosity solutions of integrodifferential equations for nonruin probabilities. Theory of Probability and its Applications, 60 (4):671–679, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bevrani:2017:SRA

[BK17] H. Bevrani and V. Yu. Korolev. Some remarks on the asymptotic behavior of the empirical availability function. Theory of Probability and its Applications, 61 (2):327–335, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bogachev:2015:TPV

[BKNP15] V. I. Bogachev, E. D. Kosov, I. Nourdin, and G. Poly. Two properties of vectors of quadratic forms in Gaussian random variables. Theory of Probability and its Applications, 59(2):208–221, ???? 2015. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bogachev:2018:DBS

[BKS18] V. I. Bogachev, A. I. Kirillov, and S. V. Shaposhnikov. Distances between stationary distributions of diffusions and solvability of nonlinear Fokker–Planck– Kolmogorov equations. Theory of Probability and its Applications, 62(1):12–34, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bickel:2012:AIB

[BL12] P. Bickel and M. Lindner. Approximating the inverse of banded matrices by banded matrices with applications to probability and statistics. Theory of Probability and its Applications, 56(1):1–20, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v56/i1/p1_s1.

Borovkov:2012:CTE

[BM12a] A. A. Borovkov and A. A. Mogulskii. Chebyshev-type exponential inequalities for sums of random vectors and for trajectories of random walks. Theory of Probability and its Applications, 56(1):21–43, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v56/i1/p21_s1.

Borovkov:2012:LDP

[BM12b] A. A. Borovkov and A. A. Mogulskii. Large deviation principles for random walk trajectories. I. Theory of Probability and its Applications, 56(4):538–561, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2013:LDPa

[BM13a] A. A. Borovkov and A. A. Mogulskii. Large deviation principles for random walk trajectories. I. Theory of Probability and its Applications, 56(4):538–561, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2013:LDPc

[BM13b] A. A. Borovkov and A. A. Mogulskii. Large deviation principles for random walk trajectories. II. Theory of Probability and its Applications, 57(1):1–27, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2014:LDP

[BM14a] A. A. Borovkov and A. A. Mogulskii. Large deviation principles for random walk trajectories. III. Theory of Probability and its Applications, 58(1):25–37, ???? 2014.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2014:MLD

[BM14b] A. A. Borovkov and A. A. Mogulskii. Moderately large deviation principles for the trajectories of random walks and processes with independent increments. Theory of Probability and its Applications, 58(4):562–581, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Boguslavskaya:2016:ESO

[BM16a] E. Boguslavskaya and D. Muravey.

An explicit solution for optimal investment in Heston model. Theory of Probability and its Applications, 60(4):679–688, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2016:LDPa

[BM16b] A. A. Borovkov and A. A. Mogulskii. Large deviation principles for trajectories of compound renewal processes. I. Theory of Probability and its Applications, 60 (2):207–224, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2016:LDPb

[BM16c] A. A. Borovkov and A. A. Mogulskii. Large deviation principles for trajectories of compound renewal processes. II. Theory of Probability and its Applications, 60 (3):349–366, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bryc:2016:WLS

[BM16d] W. Bryc and W. Matysiak. Wilson’s 6 − j laws and stitched Markov processes. Theory of Probability and its Applications, 60 (2):225–245, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borisov:2012:DNC

[BN12] I. S. Borisov and N. N. Nikitina. The distribution of the number of crossings of a strip by paths of the simplest random walks and of a Wiener process with drift. Theory of Probability and its Applications, 56(1):126–132, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v56/i1/p126_s1.

Babilua:2019:OHT

[BN19] P. K. Babilua and E. A. Nadaraya. On one homogeneity test based on quadratic deviations between kernel estimators of a distribution density in p ≥ 2 independent samples. Theory of Probability and its Applications, 63 (4):532–544, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Barmalzan:2018:ORA

[BNB18] G. Barmalzan, A. T. Payandeh Najafabadi, and N. Balakrishnan. Ordering results for aggregate claim amounts from two heterogeneous Marshall–Olkin extended exponential portfolios and their applications in insurance analysis. Theory of Probability and its Applications, 62(1):117–131, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Barndorff-Nielsen:2017:IST

[BNHS17] O. E. Barndorff-Nielsen, E. Hedevang, and J. Schmiegel. Incremental similarity and turbulence. Theory of Probability and its Applications, 61(3):482–489, ???? 2017. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Barndorff-Nielsen:2012:MTE

[BNP12] O. E. Barndorff-Nielsen and J. Pedersen. Meta-times and extended subordination. Theory of Probability and its Applications, 56(2):319–327, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bezhaeva:2013:EEM

[BO13] Z. I. Bezhaeva and V. I. Oseledets.

The entropy of the Erdős measure for the pseudogolden ratio. Theory of Probability and its Applications, 57(1):135–144, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Basse-OConnor:2014:SIR

[BOGP14] A. Basse-O’Connor, S.-E. Graversen, and J. Pedersen. Stochastic integration on the real line. Theory of Probability and its Applications, 58(2):193–215, ???? 2014. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Boldin:2013:RST

[Bol13] M. V. Boldin. Robustness of sign tests for testing hypotheses about order of autoregression. Theory of Probability and its Applications, 57(4):663–670, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Boldin:2019:RST

[Bol19] M. V. Boldin. Robust sign test for the unit root hypothesis of autoregression. Theory of Probability and its Applications, 63 (3):351–363, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2010:ILL

[Bor10] A. A. Borovkov. Integro-local and local theorems on normal and large deviations of the sums of nonidentically distributed random variables in the triangular array scheme. Theory of Probability and its Applications, 54 (4):571–587, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borisov:2012:NSL

[Bor12] I. S. Borisov. News of scientific life. Theory of Probability and its Applications, 56(4):711, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borisov:2013:NSL

[Bor13] I. S. Borisov. News of scientific life. Theory of Probability and its Applications, 56(4):711, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borisov:2014:TSN

[Bor14] I. S. Borisov. Transformations of the simplest nonsymmetric random walks and some applications of the invariance principle. Theory of Probability and its Applications, 58(2):323–329, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2015:SOA

[Bor15] A. A. Borovkov. Second order approximation for the distribution of the maximum of a random walk with negative drift and infinite variance. Theory of Probability and its Applications, 59(1):3–22, ???? 2015. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2017:GRI

[Bor17a] A. A. Borovkov. Generalization and refinement of the integro-local Stone theorem for sums of random vectors. Theory of Probability and its Applications, 61 (4):590–612, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2017:DFP

[Bor17b] A. A. Borovkov. On the distribution of the first passage time of a random walk to an arbitrary remote boundary. Theory of Probability and its Applications, 61(2):235–254, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2018:ILL

[Bor18a] A. A. Borovkov. Integro-local limit theorems for compound renewal processes. Theory of Probability and its Applications, 62 (2):175–195, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2018:EPC

[Bor18b] A. A. Borovkov. On estimation of parameters in the case of discontinuous densities. Theory of Probability and its Applications, 63 (2):169–192, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2019:SCM

[Bor19] A. A. Borovkov. Short communications moderate large deviations principles for trajectories of compound renewal processes. Theory of Probability and its Applications, 64(2):324–333, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Boyarinov:2012:PMT

[Boy12] R. N. Boyarinov. Probabilistic methods in the theory of the Riemann zeta-function. Theory of Probability and its Applications, 56 (2):181–192, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belomestny:2015:SCI

[BP15] D. V. Belomestny and A. V.

Prokhorov. Stability of characterization of the independence of random variables by the independence of linear statistics. Theory of Probability and its Applications, 59(4):672–677, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Berdjane:2016:SOC

[BP16] B. Berdjane and S. Pergamenshchikov. Sequential δ-optimal consumption and investment for stochastic volatility markets with unknown parameters. Theory of Probability and its Applications, 60 (4):533–560, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bianchi:2011:TID

[BRKF11] M. L. Bianchi, S. T. Rachev, Y. S.

Kim, and F. J. Fabozzi. Tempered infinitely divisible distributions and processes. Theory of Probability and its Applications, 55(1):2– 26, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i1/p2_s1.

Belomestny:2010:OSI

[BRU10] D. V. Belomestny, L. Rüschendorf, and M. A. Urusov. Optimal stopping of integral functionals and a “No-Loss” free boundary formulation. Theory of Probability and its Applications, 54(1):14–28, ???? 2010. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Bakirov:2011:BCC

[BS11] N. K. Bakirov and G. J. Székely. Brownian covariance and Central Limit Theorem for stationary sequences. Theory of Probability and its Applications, 55(3):371–394, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v55/i3/p371_s1.

Belomestny:2014:CIS

[BS14] D. Belomestny and V. Spokoiny. Concentration inequalities for smooth random fields. Theory of Probability and its Applications, 58(2):314–323, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulinskaya:2018:LTG

[BS18] E. V. Bulinskaya and A. I. Sokolova. Limit theorems for generalized renewal processes. Theory of Probability and its Applications, 62(1):35–54, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulinskaya:2011:CBR

[Bul11] E. Vl. Bulinskaya. Catalytic branching random walk on a two-dimensional lattice. Theory of Probability and its Applications, 55(1):120–126, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i1/p120_s1.

Bulinskaya:2015:CCC

[Bul15] E. Vl. Bulinskaya. Complete classification of catalytic branching processes. Theory of Probability and its Applications, 59 (4):545–566, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulinski:2016:EIN

[Bul16] A. V. Bulinski. Estimate of the interaction neighborhood radius for a Markov random field. Theory of Probability and its Applications, 60(2):313–318, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulinski:2017:CCL

[Bul17] A. V. Bulinski. Conditional central limit theorem. Theory of Probability and its Applications, 61 (4):613–631, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bulinskii:2018:NSL

[Bul18] A. V. Bulinskii. News of scientific life. Theory of Probability and its Applications, 63 (1):164–167, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Burkin:2016:SSR

[Bur16] A. V. Burkin. Small subgraphs in random distance graphs. Theory of Probability and its Applications, 60(3):367–382, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Butkovsky:2014:EPN

[But14] O. A. Butkovsky. On ergodic properties of nonlinear Markov chains and stochastic McKean– Vlasov equations. Theory of Probability and its Applications, 58 (4):661–674, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borisov:2016:AED

[BV16] I. S. Borisov and N. V. Volodko. Asymptotic expansions for the distributions of canonical V -statistics of third order. Theory of Probability and its Applications, 60(1):1– 18, ???? 2016. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bao:2016:CLS

[BX16] Z. Bao and J. Xie. CLT for linear spectral statistics of Hermitian Wigner matrices with general moment conditions. Theory of Probability and its Applications, 60(2):187–206, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bystrov:2015:EID

[Bys15] A. A. Bystrov. Exponential inequalities for the distribution tails of multiple stochastic integrals with respect to Gaussian integrating processes. Theory of Probability and its Applications, 59 (1):128–136, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bayraktar:2017:OSP

[BZ17] E. Bayraktar and Zh. Zhou. On an optimal stopping problem of an insider. Theory of Probability and its Applications, 61 (1):129–133, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chebotarev:2015:RFN

[CDD15] A. M. Chebotarev, V. G. Danilov, and S. Yu. Dobrokhotov. The Russian Federation National Award in Science and Technology. Theory of Probability and its Applications, 59(2):181–184, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chevallier:2011:LSB

[Che11] N. Chevallier. Law of the sum of Bernoulli random variables. Theory of Probability and its Applications, 55(1):27–41, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i1/p27_s1.

Chen:2015:SLP

[Che15] X. Chen. Scaling limit of the path leading to the leftmost particle in a branching random walk. Theory of Probability and its Applications, 59(4):567–589, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:2016:BLL

[Chi16] D. M. Chibisov. Bernoulli’s law of large numbers and the strong law of large numbers. Theory of Probability and its Applications, 60 (2):318–319, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chibisov:2019:MVV

[Chi19] D. M. Chibisov. In memory of O. V. Viskov (02.02.1937– 21.06.2018). Theory of Probability and its Applications, 63(4):671, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Czyzewska-Jankowska:2011:FPG

[CJJ11] A. Czyzewska-Jankowska and Z. J. Jurek. Factorization property of generalized s-selfdecomposable measures and class L^{f} distributions. Theory of Probability and its Applications, 55(4):692–698, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i4/p692_s1.

Chigansky:2019:SAM

[CK19] P. Chigansky and M. Kleptsyna. Statistical analysis of the mixed fractional Ornstein– Uhlenbeck process. Theory of Probability and its Applications, 63 (3):408–425, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cuchiero:2016:NPF

[CKT16] C. Cuchiero, I. Klein, and J. Teichmann. A new perspective on the fundamental theorem of asset pricing for large financial markets. Theory of Probability and its Applications, 60(4):561–579, ???? 2016. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chigansky:2010:MDD

[CL10] P. Chigansky and R. Liptser. Moderate deviations for a diffusiontype process in a random environment. Theory of Probability and its Applications, 54(1):29–50, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chobanyan:2015:MIR

[CLS15] S. Chobanyan, S. Levental, and H. Salehi. Maximum inequalities for rearrangements of summands and assignments of signs. Theory of Probability and its Applications, 59(4):677–684, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Choulli:2017:EDH

[CM17] T. Choulli and J. Ma. Explicit description of HARA forward utilities and their optimal portfolios. Theory of Probability and its Applications, 61(1):57–93, ???? 2017. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cammarota:2011:APR

[CO11] V. Cammarota and E. Orsingher. Angular processes related to Cauchy random walks. Theory of Probability and its Applications, 55(3):395–410, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v55/i3/p395_s1.

Cammarota:2013:HSH

[CO13] V. Cammarota and E. Orsingher. Hitting spheres on hyperbolic spaces. Theory of Probability and its Applications, 57(3):419–443, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Comman:2018:IPS

[Com18] H. Comman. Improvements of Plachky–Steinebach Theorem. Theory of Probability and its Applications, 63(1):117–134, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cordero:2011:SPF

[Cor11] F. Cordero. On the scaling property in fluctuation theory for stable Lévy processes. Theory of Probability and its Applications, 55(4):683–691, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v55/i4/p683_s1.

Campos:2013:MGA

[CPC13] V. S. M. Campos, A. G. C. Pereira, and J. A. Rojas Cruz. Modeling the genetic algorithm by a nonhomogeneous Markov chain: Weak and strong ergodicity. Theory of Probability and its Applications, 57 (1):144–151, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cui:2017:ITQ

[CS17] X. Cui and T. Shibata. Investment timing and quantity strategies under asymmetric information. Theory of Probability and its Applications, 61(1):151–159, ???? 2017. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cawston:2010:CPO

[CV10] S. Cawston and L. Vostrikova. On continuity properties for option prices in exponential Lévy models. Theory of Probability and its Applications, 54(4):588–608, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Darkhovsky:2012:UCP

[Dar12] B. S. Darkhovsky. Uncertain change-point problem for stochastic sequence. Theory of Probability and its Applications, 56(1):44–56, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v56/i1/p44_s1.

Darkhovsky:2014:CPD

[Dar14] B. S. Darkhovsky. Changepoint detection in random sequence under minimal prior information. Theory of Probability and its Applications, 58(3):

488–493, ????  2014.  CO-

DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dehay:2013:LTC

[Deh13] D. Dehay. Local time and convergence of empirical estimators. Theory of Probability and its Applications, 57(2):196–208, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Demichev:2014:OEC

[Dem14] V. P. Demichev. An optimal estimate for the covariance of indicator functions of associated random variables. Theory of Probability and its Applications, 58 (4):675–683, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dinev:2013:ABE

[DM13] T. Dinev and L. Mattner. The asymptotic Berry–Esseen constant for intervals. Theory of Probability and its Applications, 57 (2):323–325, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dorogovtsev:2010:NSL

[Dor10] A. A. Dorogovtsev. News of scientific life: International conference “Stochastic Analysis and Random Dynamical Systems”. Theory of Probability and its Applications, 54(4):733, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davydov:2013:SAL

[DP13] Yu. Davydov and V. Paulauskas.

A simple approach in limit theorems for linear random processes and fields with continuous time. Theory of Probability and its Applications, 57(4): 589–606, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Doumas:2018:SCC

[DP18a] A. V. Doumas and V. G. Papanicolaou. The siblings of the coupon collector. Theory of Probability and its Applications, 62 (3):444–470, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

DOvidio:2018:FDT

[DP18b] M. D’Ovidio and F. Polito. Fractional Diffusion–Telegraph equations and their associated stochastic solutions. Theory of Probability and its Applications, 62 (4):552–574, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davydov:2019:LTT

[DP19] Yu. Davydov and V. Paulauskas.

Lamperti-type theorems for random fields. Theory of Probability and its Applications, 63 (3):426–446, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

deChaumaray:2017:LDS

[dRdC17] M. du Roy de Chaumaray. Large deviations for the squared radial Ornstein–Uhlenbeck process. Theory of Probability and its Applications, 61(3):408–441, ???? 2017. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Denisov:2019:FPT

[DSW19] D. Denisov, A. Sakhanenko, and V. Wachtel. First-passage times over moving boundaries for asymptotically stable walks. Theory of Probability and its Applications, 63(4):613–633, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dyakonova:2012:MGW

[Dya12] E. E. Dyakonova. Multitype Galton–Watson branching processes in Markovian random environment. Theory of Probability and its Applications, 56 (3):508–517, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Embrechts:2018:HG

[EK18] P. Embrechts and M. Kirchner. Hawkes graphs. Theory of Probability and its Applications, 62 (1):132–156, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Engelbert:2019:TFS

[EK19] H.-J. Engelbert and V. P. Kurenok. The Tanaka formula for symmetric stable processes with index α, 0 < α < 2. Theory of Probability and its Applications, 64(2):264–289, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Es-Sebaiy:2011:NLT

[EST11] K. Es-Sebaiy and C. A. Tudor. Noncentral limit theorem for the cubic variation of a class of self-similar stochastic processes. Theory of Probability and its Applications, 55(3):411–431, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i3/p411_s1.

Eliseeva:2013:ECF

[EZ13] Yu. S. Eliseeva and A. Yu. Zaitsev.

Estimates of the concentration functions of weighted sums of independent random variables. Theory of Probability and its Applications, 57(4):670–678, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fatalov:2014:LMG

[Fat14] V. R. Fatalov. On the Laplace method for Gaussian measures in a Banach space. Theory of Probability and its Applications, 58 (2):216–241, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fatalov:2015:LE

[Fat15] V. R. Fatalov. Letter to the Editors. Theory of Probability and its Applications, 59(2):357–358, ???? 2015. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fazekas:2012:MSP

[Faz12] I. Fazekas. Merging to semistable processes. Theory of Probability and its Applications, 56(4): 621–633, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fazekas:2013:MSP

[Faz13] I. Fazekas. Merging to semistable processes. Theory of Probability and its Applications, 56(4): 621–633, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feldman:2010:ILS

[Fel10] G. M. Feldman. Independent linear statistics on a-adic solenoids. Theory of Probability and its Applications, 54(3):375–388, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v54/i3/p375_s1.

Feldman:2012:GMS

[Fel12] G. M. Feldman. Gaussian measures in the sense of Bernstein: Factorization, supports, zero-one law. Theory of Probability and its Applications, 56(3):359–375, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feldman:2017:IRV

[Fel17] G. M. Feldman. Independent random variables in Abelian groups with independent sum and difference. Theory of Probability and its Applications, 61(2):335–345, ???? 2017. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feldman:2018:HCT

[Fel18] G. M. Feldman. Heyde’s characterization theorem for some locally compact Abelian groups. Theory of Probability and its Applications, 62(3):399–412, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feldman:2019:CTP

[Fel19] G. M. Feldman. On a characterization theorem for probability distributions on discrete Abelian groups. Theory of Probability and its Applications, 63(4):594–612, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feinberg:2014:FLW

[FKZ14] E. A. Feinberg, P. O. Kasyanov, and N. V. Zadoianchuk. Fatou’s lemma for weakly converging probabilities. Theory of Probability and its Applications, 58 (4):683–689, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feinberg:2010:SEN

[FP10] E. A. Feinberg and A. B. Piunovskiy. On strongly equivalent nonrandomized transition probabilities. Theory of Probability and its Applications, 54(2):300–307, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v54/i2/p300_s1.

Faller:2013:OMS

[FR13] A. Faller and L. Rüschendorf. Optimal multiple stopping with sum-payoff. Theory of Probability and its Applications, 57 (2):325–336, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ferger:2010:WCE

[FV10] D. Ferger and D. Vogel. Weak convergence of the empirical process and the rescaled empirical distribution function in the Skorokhod product space. Theory of Probability and its Applications, 54 (4):609–625, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gamkrelidze:2015:LLT

[Gam15] N. G. Gamkrelidze. On a local limit theorem for integer random vectors. Theory of Probability and its Applications, 59 (3):494–499, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gamkrelidze:2018:OPS

[Gam18] N. G. Gamkrelidze. On one property of symmetrized distributions. Theory of Probability and its Applications, 62(1):55–57, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gaposhkin:2011:ECR

[Gap11] V. F. Gaposhkin. On estimates of the convergence rate in the strong law of large numbers for stationary sequences (in some classes S_{w} and R_{Φ}). Theory of Probability and its Applications, 55(4):699–709, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i4/p699_s1.

Gusev:2015:DFC

[GD15] S. A. Gusev and N. G. Dokuchaev.

On differentiation of functionals containing the first exit of a diffusion process from a domain. Theory of Probability and its Applications, 59(1):136–144, ???? 2015. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gotze:2018:AIC

[GEZ18] F. Götze, Yu. S. Eliseeva, and A. Yu. Zaitsev. Arak inequalities for concentration functions and the Littlewood–Offord problem. Theory of Probability and its Applications, 62(2):196–215, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gao:2014:BTC

[GGW14] F. Gao, A. Guillin, and L. Wu.

Bernstein-type concentration inequalities for symmetric Markov processes. Theory of Probability and its Applications, 58(3):

358–382, ????  2014.  CO-

DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gribkova:2011:CBA

[GH11] N. V. Gribkova and R. Helmers. On the consistency of the M ≪ N bootstrap approximation for a trimmed mean. Theory of Probability and its Applications, 55(1):42–53, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v55/i1/p42_s1.

Gribkova:2014:SOA

[GH14] N. V. Gribkova and R. Helmers. Second order approximations for slightly trimmed means. Theory of Probability and its Applications, 58(3):383–412, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gyorfi:2015:ABM

[GIKP15] L. Györfi, M. Ispány, P. Kevei, and G. Pap. Asymptotic behavior of multitype nearly critical Galton–Watson processes with immigration. Theory of Probability and its Applications, 59 (4):590–610, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Goldenshluger:2013:GSR

[GL13] A. V. Goldenshluger and O. V. Lepski. General selection rule from a family of linear estimators. Theory of Probability and its Applications, 57(2):209–226, ???? 2013.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Glau:2016:CLP

[Gla16] K. Glau. Classification of Lévy processes with parabolic Kolmogorov backward equations. Theory of Probability and its Applications, 60(3):383–406, ???? 2016. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grabchak:2015:LTP

[GM15] M. Grabchak and S. A. Molchanov.

Limit theorems and phase transitions for two models of summation of independent identically distributed random variables with a parameter. Theory of Probability and its Applications, 59 (2):222–243, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grabchak:2019:LTR

[GM19] M. Grabchak and S. A. Molchanov.

Limit theorems for random exponentials: The bounded support case. Theory of Probability and its Applications, 63(4):634–647, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gotze:2015:LTT

[GNT15] F. Götze, A. A. Naumov, and A. N. Tikhomirov. Limit theorems for two classes of random matrices with dependent entries. Theory of Probability and its Applications, 59(1):23–39, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gotze:2018:LSL

[GNT18] F. Götze, A. A. Naumov, and A. N. Tikhomirov. Local semicircle law under moment conditions: The Stieltjes transform, rigidity, and delocalization. Theory of Probability and its Applications, 62(1):58–83, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Golosov:2013:ULS

[Gol13] A. O. Golosov. Uniform laws for stationary random functions on compact groups. Theory of Probability and its Applications, 57 (2):336–342, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Goldaeva:2014:EIC

[Gol14] A. A. Goldaeva. Extremal indices and clusters in the linear recursive stochastic sequences. Theory of Probability and its Applications, 58(4):689–698, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grbic:2010:GPT

[GP10] T. Grbić and E. Pap. Generalization of portmanteau theorem with respect to the pseudoweak convergence of random closed sets. Theory of Probability and its Applications, 54(1):51–67, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Glonti:2017:OIR

[GP17] O. A. Glonti and O. G. Purtukhia. On one integral representation of functionals of Brownian motion. Theory of Probability and its Applications, 61 (1):133–139, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gregoratti:2010:DQP

[Gre10] M. Gregoratti. Dilations à la quantum probability of Markov evolutions in discrete time. Theory of Probability and its Applications, 54(1):140–150, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grin:2010:MCW

[Gri10] A. G. Grin’. On minimal conditions of the weak dependence in limit theorems for stationary sequences. Theory of Probability and its Applications, 54(2):307–317, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i2/p307_s1.

Grishunina:2019:MQS

[Gri19] S. A. Grishunina. Multiserver queueing system with constant service time and simultaneous service of a customer by random number of servers. Theory of Probability and its Applications, 64 (3):456–460, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ginovyan:2012:TAP

[GS12] M. S. Ginovyan and A. A. Sahakyan. Trace approximations of products of truncated Toeplitz operators. Theory of Probability and its Applications, 56(1):57–71, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v56/i1/p57_s1.

Gliklikh:2019:CSF

[GS19] Yu. E. Gliklikh and T. A. Shchichko. On the completeness of stochastic flows generated by equations with current velocities. Theory of Probability and its Applications, 64(1):1–11, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gotze:2010:RCS

[GT10] F. Götze and A. N. Tikhomirov. The rate of convergence of spectra of sample covariance matrices. Theory of Probability and its Applications, 54(1):129–140, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gushchin:2016:PCE

[GU16] A. A. Gushchin and M. A. Urusov. Processes that can be embedded in a geometric Brownian motion. Theory of Probability and its Applications, 60(2):246–262, ???? 2016. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gushchin:2011:DCV

[Gus11] A. A. Gushchin. Dual characterization of the value function in the robust utility maximization problem. Theory of Probability and its Applications, 55(4):611–630, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i4/p611_s1.

Gushchin:2018:JLT

[Gus18] A. A. Gushchin. The joint law of terminal values of a nonnegative submartingale and its compensator. Theory of Probability and its Applications, 62(2): 216–235, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gotze:2013:MTY

[GvZ13] F. Götze and W. R. van Zwet.

Miscellanea — the times of Yuri Vasilyevich Prokhorov. Theory of Probability and its Applications, 57 (2):306–322, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gotze:2012:DCR

[GZ12] F. Götze and D. Zaporozhets.

On the distribution of complex roots of random polynomials with heavy-tailed coefficients. Theory of Probability and its Applications, 56(4):696–703, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gotze:2013:DCR

[GZ13] F. Götze and D. Zaporozhets. On the distribution of complex roots of random polynomials with heavy-tailed coefficients. Theory of Probability and its Applications, 56(4):696–703, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Harlamov:2016:FDD

[Har16] B. P. Harlamov. Final distribution of a diffusion process: Semi-Markov approach. Theory of Probability and its Applications, 60 (3):444–459, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Husler:2016:STG

[HKP16] J. Hüsler, E. V. Kremena, and V. I. Piterbarg. On the shape of trajectories of Gaussian processes having large massive excursions. II. Theory of Probability and its Applications, 60 (3):513–520, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hajiyev:2012:MMM

[HM12] A. H. Hajiyev and T. Sh. Mammadov. Mathematical models of moving particles without overtake and their applications. Theory of Probability and its Applications, 56(4):579–589, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hajiyev:2013:MMM

[HM13] A. H. Hajiyev and T. Sh. Mammadov. Mathematical models of moving particles without overtake and their applications. Theory of Probability and its Applications, 56(4):579–589, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hamadene:2016:RBS

[HO16] S. Hamadène and Y. Ouknine. Reflected backward SDEs with general jumps. Theory of Probability and its Applications, 60 (2):263–280, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Holevo:2016:SDM

[Hol16] A. S. Holevo. Seminar dedicated to the memory of Yu. V. Prokhorov. Theory of Probability and its Applications, 60(2):312, ???? 2016. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Holevo:2017:CGT

[Hol17] A. S. Holevo. On a characterization of Gaussian transition operators. Theory of Probability and its Applications, 61(4): 712–718, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Holevo:2018:CCC

[Hol18] A. S. Holevo. On the classical capacity of a channel with stationary quantum Gaussian noise. Theory of Probability and its Applications, 62(4):534–551, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Holevo:2019:GQO

[Hol19] A. S. Holevo. Generators of quantum one-dimensional diffusions. Theory of Probability and its Applications, 64(2):249–263, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Holevo:2016:CWC

[HS16] A. S. Holevo and M. E. Shirokov. Criterion of weak compactness for families of generalized ensembles and its corollaries. Theory of Probability and its Applications, 60 (2):320–325, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hu:2012:BRW

[HTV12] Y. Hu, V. A. Topchii, and V. A. Vatutin. Branching random walk in Z^{4} with branching at the origin only. Theory of Probability and its Applications, 56(2): 193–212, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hinz:2016:AOC

[HY16] J. Hinz and N. Yap. Algorithms for optimal control of stochastic switching systems. Theory of Probability and its Applications, 60 (4):580–603, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ham:2010:SNV

[HZ10] W. Ham and K. Zhou. A short note on the volume of hypersphere. Theory of Probability and its Applications, 54(4):682–686, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hwang:2011:UAP

[HZ11] H.-K. Hwang and V. Zacharovas.

Uniform asymptotics of Poisson approximation to the Poissonbinomial distribution. Theory of Probability and its Applications, 55(2):198–224, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i2/p198_s1.

Ibragimov:2012:MAV

[Ibr12]  I. A. Ibragimov.  In mem-

ory of Anatolii Vladimirovich Skorokhod (1930–2011). Theory of Probability and its Applications, 56(1):116–119, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v56/i1/p116_s1.

Ibragimov:2013:SDR

[Ibr13] I. A. Ibragimov. On the Skitovich– Darmois–Ramachandran theorem. Theory of Probability and its Applications, 57(3):368–374, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:2018:I

[IKS18] I. A. Ibragimov, V. Yu. Korolev, and A. N. Shiryaev. Introduction.

Theory of Probability and its Applications, 62(1):1, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ilinskii:2010:GCF

[Il’10] A. I. Il’inskii. On Gaussian components of finite mixtures of probability distributions. Theory of Probability and its Applications, 54 (4):687–693, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Iksanov:2015:WCF

[IMV15] A. M. Iksanov, A. V. Marynych, and V. A. Vatutin. Weak convergence of finite-dimensional distributions of the number of empty boxes in the Bernoulli sieve. Theory of Probability and its Applications, 59(1):87–113, ???? 2015. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Irkhina:2011:OSR

[Irk11] N. A. Irkhina. On one sufficient reducibility criterion for Wang’s premium principle. Theory of Probability and its Applications, 55(1):126–134, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i1/p126_s1.

Ivanov:2012:DPH

[IS12] R. V. Ivanov and A. N. Shiryaev. On duality principle for hedging strategies in diffusion models. Theory of Probability and its Applications, 56(3):376–402, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:2014:PMS

[ISF14] I. A. Ibragimov, N. V. Smorodina, and M. M. Faddeev. On a probabilistic method of solving a one-dimensional initial-boundary value problem. Theory of Probability and its Applications, 58 (2):242–263, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:2015:LTC

[ISF15] I. A. Ibragimov, N. V. Smorodina, and M. M. Faddeev. Limit theorems on convergence of expectations of functionals of sums of independent random variables to solutions of initial-boundary value problems. Theory of Probability and its Applications, 59 (2):244–259, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:2017:ADP

[ISF17a] I. A. Ibragimov, N. V. Smorodina, and M. M. Faddeev. Analytic diffusion processes: Definition, properties, limit theorems. Theory of Probability and its Applications, 61 (2):255–276, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:2017:IBV

[ISF17b] I. A. Ibragimov, N. V. Smorodina, and M. M Faddeev. Initial boundary value problems in a bounded domain: Probabilistic representations of solutions and limit theorems. I. Theory of Probability and its Applications, 61 (4):632–648, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:2018:IBV

[ISF18] I. A. Ibragimov, N. V. Smorodina, and M. M. Faddeev. Initial boundary value problems in a bounded domain: Probabilistic representations of solutions and limit theorems II. Theory of Probability and its Applications, 62 (3):356–372, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:2019:AEO

[ISF19a] I. A. Ibragimov, N. V. Smorodina, and M. M. Faddeev. Approximation of the evolution operator by expectations of functionals of sums of independent random variables. Theory of Probability and its Applications, 64(1):12– 26, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:2019:RLP

[ISF19b] I. A. Ibragimov, N. V. Smorodina, and M. M. Faddeev. Reflecting Lévy processes and associated families of linear operators. Theory of Probability and its Applications, 64(3):335–354, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivanov:2015:CBP

[Iva15] M. Yu. Ivanov. On a connection between power and logarithmic utility maximization problems in the exponential Lévy model. Theory of Probability and its Applications, 59(4):684–693, ???? 2015. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Jajte:2010:OSD

[Jaj10] R. Jajte. Operators on L^{p}- spaces determined by filtrations and pointwise ergodic theorems. Theory of Probability and its Applications, 54(3):389–402, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i3/p389_s1.

Jasiulis-Goldyn:2016:WLK

[JGM16] B. H. Jasiulis-Goldyn and J. K.

Misiewicz. Weak Lévy–Khintchine representation for weak infinite divisibility. Theory of Probability and its Applications, 60(1):45–61, ???? 2016. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Jaimungal:2014:GSP

[JKV14] S. Jaimungal, A. Kreinin, and A. Valov. The generalized Shiryaev problem and Skorokhod embedding. Theory of Probability and its Applications, 58 (3):493–502, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Jeanblanc:2015:MMY

[JS15] M. Jeanblanc and A. Shiryaev. In memory of Marc Yor. Theory of Probability and its Applications, 59(1):180, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Jana:2016:FLE

[JSS16] I. Jana, K. Saha, and A. Soshnikov. Fluctuations of linear eigenvalue statistics of random band matrices. Theory of Probability and its Applications, 60 (3):407–443, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kaji:2017:FPP

[Kaj17] Sh. Kaji. First passage problems over increasing boundaries for Lévy processes with exponentially decayed Lévy measures. Theory of Probability and its Applications, 61(1):140–151, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kareev:2013:LBA

[Kar13] I. A. Kareev. Lower bounds for average sample size and efficiency of sequential selection procedures. Theory of Probability and its Applications, 57(2):227–242, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kareev:2014:LBA

[Kar14] I. A. Kareev. Lower bound for the average sample size and the efficiency of ranking sequential procedures. Theory of Probability and its Applications, 58 (3):503–509, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Karpushin:2017:CRS

[Kar17] V. V. Karpushin. Convergence rate in stochastic particle systems with synchronization. Theory of Probability and its Applications, 61 (2):345–355, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kashitsyn:2012:MMC

[Kas12] P. A. Kashitsyn. Multivariate model with correlated observation units. Theory of Probability and its Applications, 56 (3):517–521, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kashtanov:2016:SFA

[Kas16] V. A. Kashtanov. The structure of the functional of accumulation defined on a trajectory of semi-Markov process with a finite set of states. Theory of Probability and its Applications, 60 (2):281–294, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korolev:2018:BCF

[KD18] V. Yu. Korolev and A. V. Dorofeeva. Bounds for the concentration functions of random sums under relaxed moment conditions. Theory of Probability and its Applications, 62(1):84–97, ???? 2018.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kelbert:2010:MPR

[Kel10] M. Kelbert. Markov process representations for polyharmonic functions. Theory of Probability and its Applications, 54(2):317–322, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v54/i2/p317_s1.

Kern:2016:GMV

[Ker16] P. Kern. A general multiparameter version of Gnedenko’s transfer theorem. Theory of Probability and its Applications, 60 (1):134–142, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuznetsova:2014:CTH

[KH14] A. A. Kuznetsova and A. S. Holevo. Coding theorems for hybrid channels. Theory of Probability and its Applications, 58 (2):264–285, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuznetsova:2015:CTH

[KH15] A. A. Kuznetsova and A. S. Holevo. Coding theorems for hybrid channels. II. Theory of Probability and its Applications, 59 (1):145–154, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2010:BS

[Kha10] Marina Khatuntseva. On the 75th birthday of A. N. Shiryaev. Theory of Probability and its Applications, 54(4):677, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2012:C

[Kha12a] M. V. Khatuntseva. Congratulations. Theory of Probability and its Applications, 56(2):318, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2012:NSL

[Kha12b] M. V. Khatuntseva. News of scientific life. Theory of Probability and its Applications, 56(3):536, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2012:BB

[Kha12c] M. V. Khatuntseva. On the 80th birthday of A. A. Borovkov. Theory of Probability and its Applications, 56(4):537, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khasanov:2013:UHP

[Kha13a] R. V. Khasanov. On the upper hedging price of contingent claims. Theory of Probability and its Applications, 57(4):607–618, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2013:NSL

[Kha13b] M. V. Khatuntseva. News of scientific life. Theory of Probability and its Applications, 57 (4):721–724, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2013:BVV

[Kha13c] M. V. Khatuntseva. On the 60th birthday of V. A. Vatutin. Theory of Probability and its Applications, 57(1):181, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2013:BB

[Kha13d] M. V. Khatuntseva. On the 80th birthday of A. A. Borovkov. Theory of Probability and its Applications, 56(4):537, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2014:BAS

[Kha14a] M. Khatuntseva. On the 70th birthday of Aleksander Semenovich Holevo. Theory of Probability and its Applications, 58(4):660, ???? 2014. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2014:MBS

[Kha14b] M. V. Khatuntseva. In memoriam: B. A. Sevastyanov (29.09.1923– 30.08.2013). Theory of Probability and its Applications, 58(3):487, ???? 2014. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2014:MYV

[Kha14c] M. V. Khatuntseva. In memoriam:

Yu. V. Prokhorov (15.12.1929– 16.07.2013). Theory of Probability and its Applications, 58(2):173, ???? 2014. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2015:MOB

[Kha15] M. V. Khatuntseva. In memoriam: [obituaries of Bronius Grigelionis, Nikolai Vakhania, and Aleksei Chervonenkis]. Theory of Probability and its Applications, 59 (3):532–534, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2016:CSA

[Kha16a] M. V. Khatuntseva. The Claude E. Shannon Award. Theory of Probability and its Applications, 60(3):501, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2016:CS

[Kha16b] M. V. Khatuntseva. Congratulations to A. N. Shiryaev. Theory of Probability and its Applications, 60(4):670, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2016:E

[Kha16c] M. V. Khatuntseva. From the Editors. Theory of Probability and its Applications, 60(3):529, ???? 2016. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2017:NSL

[Kha17a] M. V. Khatuntseva. News of scientific life. Theory of Probability and its Applications, 61(2):356, ???? 2017. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2017:BB

[Kha17b] M. V. Khatuntseva. On the 85th birthday of A. A. Borovkov. Theory of Probability and its Applications, 61(2):177, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khartov:2018:CRS

[Kha18a] A. A. Khartov. Criteria of relative and stochastic compactness for distributions of sums of independent random variables. Theory of Probability and its Applications, 63(1):57–71, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2018:NSL

[Kha18b] M. V. Khatuntseva. News of scientific life. Theory of Probability and its Applications, 63(2):336, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khikhol:2010:FGD

[Khi10] S. A. Khikhol. A formula for the generalized density process of the distributions of semimartingales with independent increments. Theory of Probability and its Applications, 54(4):

626–637, ????  2010.  CO-

DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khikhol:2014:ALC

[Khi14a] S. A. Khikhol. Averaging the local characteristics brings a semimartingale with independent increments closer to Lévy processes. Theory of Probability and its Applications, 58(3):413–429, ???? 2014. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khil:2014:MSS

[Khi14b] E. V. Khil. Markov structure of the sequence of local maxima and the gaps between them. Theory of Probability and its Applications, 58(3):430–441, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khokhlov:2013:RB

[Kho13] Yu. S. Khokhlov. Reviews and bibliography. Theory of Probability and its Applications, 57 (4):718–720, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klebaner:2011:AAR

[KL11] F. Klebaner and R. Liptser.

Asymptotic analysis of ruin in the constant elasticity of variance model. Theory of Probability and its Applications, 55(2):291–297, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i2/p291_s1.

Klebaner:2014:WSE

[KL14] F. Klebaner and R. Liptser. When a stochastic exponential is a true martingale. Extension of the Benes method. Theory of Probability and its Applications, 58(1):38–62, ???? 2014. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kabanov:2018:CFN

[KME18] Yu. M. Kabanov, R. Mokbel, and Kh. El Bitar. Clearing in financial networks. Theory of Probability and its Applications, 62 (2):252–277, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kagan:2015:PCS

[KMM15] A. M. Kagan, Y. Malinovsky, and L. Mattner. Partially complete sufficient statistics are jointly complete. Theory of Probability and its Applications, 59(3):359–374, ???? 2015. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korolyuk:2014:NSL

[KMS14] V. S. Korolyuk, Yu. S. Mishura, and L. M. Sakhno. News of scientific life. Theory of Probability and its Applications, 58(1):172, ???? 2014. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kolodii:2012:PMS

[Kol12] N. A. Kolodii. On parametermeasurability of the stochastic integral with respect to the two-parameter strong martingale. Theory of Probability and its Applications, 56(1):132–140, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v56/i1/p132_s1.

Kolesnikov:2013:SRM

[Kol13] A. V. Kolesnikov. On Sobolev regularity of mass transport and transportation inequalities. Theory of Probability and its Applications, 57(2):243–264, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Konovalov:2010:CRM

[Kon10] L. A. Konovalov. Coherent risk measures and a limit pass. Theory of Probability and its Applications, 54(3):403–423, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i3/p403_s1.

Konarovskii:2011:ISD

[Kon11a] V. V. Konarovskii. On infinite system of diffusing particles with coalescing. Theory of Probability and its Applications, 55(1):134–144, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i1/p134_s1.

Konovalov:2011:OLR

[Kon11b] L. A. Konovalov. On one limit relation for coherent risk measures. Theory of Probability and its Applications, 55(1):144–153, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v55/i1/p144_s1.

Korolev:2014:GHL

[Kor14] V. Yu. Korolev. Generalized hyperbolic laws as limit distributions for random sums. Theory of Probability and its Applications, 58(1):63–75, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korolev:2017:LDD

[Kor17] V. Yu. Korolev. Limit distributions for doubly stochastically rarefied renewal processes and their properties. Theory of Probability and its Applications, 61(4): 649–664, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kovalenko:2014:LWB

[Kov14] I. N. Kovalenko. The life and works of Boris Vladimirovich Gnedenko: a survey. Theory of Probability and its Applications, 58(1):4–16, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kozlov:2010:LDS

[Koz10] M. V. Kozlov. On large deviations of strictly subcritical branching processes in a random environment with geometric distribution of progeny. Theory of Probability and its Applications, 54(3):424–446, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i3/p424_s1.

Kozlov:2014:LDM

[Koz14] M. V. Kozlov. On large deviations of maximum of a Cramér random walk and the queueing process. Theory of Probability and its Applications, 58(1):76–106, ???? 2014. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kozhina:2017:SDP

[Koz17] A. A. Kozhina. Stability of densities for perturbed degenerate diffusions. Theory of Probability and its Applications, 61 (3):489–499, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korolev:2012:ICR

[KP12] V. Yu. Korolev and S. V. Popov.

An improvement of convergence rate estimates in the Central Limit Theorem under absence of moments higher than the second. Theory of Probability and its Applications, 56(4):682–691, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korolev:2013:ICR

[KP13] V. Yu. Korolev and S. V. Popov. An improvement of convergence rate estimates in the Central Limit Theorem under absence of moments higher than the second. Theory of Probability and its Applications, 56(4):682–691, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kleban:2019:MME

[KP19] A. O. Kleban and V. I. Piterbarg. Method of moments for exit probabilities of Gaussian vector processes from a large region. Theory of Probability and its Applications, 63(4):545–555, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kremena:2014:STG

[KPH14] E. V. Kremena, V. I. Piterbarg, and J. Hüsler. On the shape of trajectories of Gaussian processes having large massive excursions. Theory of Probability and its Applications, 58(4): 582–600, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Konev:2014:ERP

[KPP14] V. V. Konev, S. M. Pergamenshchikov, and E. A. Pchelintsev. Estimation of a regression with the pulse type noise from discrete data. Theory of Probability and its Applications, 58 (3):442–457, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kurushima:2018:NTP

[KPZ18] A. Kurushima, A. Piunovskiy, and Y. Zhang. Nowak’s theorem on probability measures induced by strategies revisited. Theory of Probability and its Applications, 62 (2):328–334, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kramkov:2016:EUA

[Kra16] D. Kramkov. Existence and uniqueness of Arrow–Debreu equilibria with consumptions in L_{+}^{0}. Theory of Probability and its Applications, 60(4):688–695, ???? 2016. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:2010:CNF

[Kru10] V. M. Kruglov. On the continuity of natural filtrations of processes with independent increments. Theory of Probability and its Applications, 54(4):693–699, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:2012:SPU

[Kru12] V. M. Kruglov. A simple proof of the unboundedness of homogeneous stochastic processes with independent increments. Theory of Probability and its Applications, 56(4):691–693, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:2013:SPU

[Kru13] V. M. Kruglov. A simple proof of the unboundedness of homogeneous stochastic processes with independent increments. Theory of Probability and its Applications, 56(4):691–693, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:2014:OID

[Kru14] V. M. Kruglov. On one identity for distribution of sums of independent random variables. Theory of Probability and its Applications, 58(2):329–331, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kryazhimskiy:2016:PIP

[Kry16] A. V. Kryazhimskiy. A posteriori integration of probabilities. Elementary theory. Theory of Probability and its Applications, 60(1):62–87, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Krylova:2019:SET

[Kry19] G. A. Krylova. Statistical estimate of the traffic coefficient for a multichannel queueing system with regenerative input flow. Theory of Probability and its Applications, 64 (3):461–464, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korolev:2010:UBA

[KS10] V. Yu. Korolev and I. G. Shevtsova. On the upper bound for the absolute constant in the Berry–Esseen inequality. Theory of Probability and its Applications, 54(4):638–658, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korolev:2011:NMT

[KS11] V. Yu. Korolev and I. G.

Shevtsova. A new momenttype estimate of convergence rate in the Lyapunov theorem. Theory of Probability and its Applications, 55(3):505–509, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i3/p505_s1.

Klartag:2012:VBE

[KS12] B. Klartag and S. Sodin. Variations on the Berry–Esseen theorem. Theory of Probability and its Applications, 56(3):403–419, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kovalevskii:2015:ASR

[KS15] A. P. Kovalevskii and E. V.

Shatalin. Asymptotics of sums of residuals of one-parameter linear regression on order statistics. Theory of Probability and its Applications, 59(3):375–387, ???? 2015. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kabanov:2016:MPF

[KS16] Y. Kabanov and A. Shiryaev.

Modern problems of financial mathematics. Theory of Probability and its Applications, 60 (4):531–532, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kabanov:2017:MPF

[KS17] Yu. M. Kabanov and A. N. Shiryaev. Modern problems of financial mathematics. Theory of Probability and its Applications, 61(1):1–2, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kampf:2018:FCL

[KS18] J. Kampf and E. Spodarev. A functional central limit theorem for integrals of stationary mixing random fields. Theory of Probability and its Applications, 63 (1):135–150, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kvaratskhelia:2014:DCR

[KT14] V. V. Kvaratskhelia and V. I. Tarieladze. Diagonally canonical and related Gaussian random elements. Theory of Probability and its Applications, 58 (2):286–296, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kudryavtsev:2019:AWH

[Kud19] O. E. Kudryavtsev. Approximate Wiener–Hopf factorization and Monte Carlo methods for Lévy processes. Theory of Probability and its Applications, 64 (2):186–208, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuznetsov:2011:APP

[Kuz11a] A. Kuznetsov. Analytic proof of Pecherskii–Rogozin identity and Wiener–Hopf factorization. Theory of Probability and its Applications, 55(3):432–443, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i3/p432_s1.

Kuznetsova:2011:CEI

[Kuz11b] A. A. Kuznetsova. Conditional entropy for infinitedimensional quantum systems. Theory of Probability and its Applications, 55(4):709–717, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i4/p709_s1.

Kuznetsova:2011:RLT

[Kuz11c] T. V. Kuznetsova. Refinement of the limit theorem for maxima of independent random sums in the case of zero asymmetry. Theory of Probability and its Applications, 55(2):298–302, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i2/p298_s1.

Klokov:2013:LMC

[KV13] S. A. Klokov and A. Yu. Veretennikov. On local mixing conditions for SDE approximations. Theory of Probability and its Applications, 57(1):110–131, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khristolyubov:2019:LTS

[KY19] I. I. Khristolyubov and E. B. Yarovaya. A limit theorem for supercritical random branching walks with branching sources of varying intensity. Theory of Probability and its Applications, 64 (3):365–384, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

LaGatta:2013:CDG

[LaG13a] T. LaGatta. Continuous disintegrations of Gaussian processes. Theory of Probability and its Applications, 57(1):151–162, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lagutin:2013:SMG

[Lag13b] M. B. Lagutin. Some model of gas penetration into porous medium. Theory of Probability and its Applications, 57(2): 342–346, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Laurent:2011:VCR

[Lau11] S. Laurent. On Vershikian and I-cosy random variables and filtrations. Theory of Probability and its Applications, 55(1):54–76, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v55/i1/p54_s1.

Linke:2018:CEE

[LB18] Yu. Yu. Linke and I. S. Borisov.

Constructing explicit estimators in nonlinear regression problems. Theory of Probability and its Applications, 63(1):22–44, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lebedev:2010:TAS

[Leb10] A. V. Lebedev. Tail asymptotics for stationary distributions of maximal branching processes. Theory of Probability and its Applications, 54(4):699–702, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lebedev:2013:MER

[Leb13] A. V. Lebedev. Multivariate extremes of random properties of particles in supercritical branching processes. Theory of Probability and its Applications, 57 (4):678–683, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lebedev:2019:UBE

[Leb19] A. V. Lebedev. Upper bound for the expected minimum of dependent random variables with known Kendall’s Tau. Theory of Probability and its Applications, 64 (3):465–473, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lopez-Garcia:2011:CSL

[LG11] M. López-García. Characterization of solutions to the lognormal moment problem. Theory of Probability and its Applications, 55(2):303–307, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i2/p303_s1.

Li:2015:LFS

[Li15] L. Li. From Lyapunov functions to Sobolev inequalities. Theory of Probability and its Applications, 59(4):693–699, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linke:2016:RFO

[Lin16] Yu. Yu. Linke. Refinement of Fisher’s one-step estimators in the case of slowly converging initial estimators. Theory of Probability and its Applications, 60(1):88–102, ???? 2016. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linke:2018:APO

[Lin18] Yu. Yu. Linke. Asymptotic properties of one-step weighted Mestimators with applications to regression. Theory of Probability and its Applications, 62(3): 373–398, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lisovskii:2019:BST

[Lis19] D. I. Lisovskii. Bayesian sequential testing problem for a Brownian bridge. Theory of Probability and its Applications, 63 (4):556–579, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lykov:2013:LHS

[LMM13] A. A. Lykov, V. A. Malyshev, and S. A. Muzychka. Linear Hamiltonian systems under microscopic random influence. Theory of Probability and its Applications, 57 (4):684–688, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lotov:2015:ARP

[Lot15] V. I. Lotov. On the asymptotics of the ruin probability. Theory of Probability and its Applications, 59(1):154–163, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lototsky:2010:UAS

[LR10] S. V. Lototsky and B. L. Rozovskii. A unified approach to stochastic evolution equations using the Skorokhod integral. Theory of Probability and its Applications, 54(2):189–202, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i2/p189_s1.

Lisovskii:2019:STT

[LS19] D. I. Lisovskii and A. N. Shiryaev.

Sequential testing of two hypotheses for a stationary Ornstein– Uhlenbeck process. Theory of Probability and its Applications, 63 (4):580–593, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Liu:2019:RCB

[LV19] M. Liu and V. Vatutin. Reduced critical branching processes for small populations. Theory of Probability and its Applications, 63 (4):648–656, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lykov:2018:RVF

[Lyk18] K. V. Lykov. Any random variable with finite moments is a sum of two variables with determinate moment problem. Theory of Probability and its Applications, 62 (4):632–639, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lyulko:2010:SRM

[Lyu10] Ya. A. Lyulko. Stochastic representations of max-type functionals of a random walk. Theory of Probability and its Applications, 54(3):516–525, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i3/p516_s1.

Lyulko:2012:DTS

[Lyu12] Ya. A. Lyulko. On the distribution of time spent by a Markov chain at different levels until achieving a fixed state. Theory of Probability and its Applications, 56(1):140–149, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v56/i1/p140_s1.

Ma:2013:SLK

[Ma13] C. Ma. The Schoenberg–Lévy kernel and relationships among fractional Brownian motion, bifractional Brownian motion, and others. Theory of Probability and its Applications, 57(4):619–632, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ma:2017:BVR

[Ma17] C. Ma. Binomial-χ^{2} vector random fields. Theory of Probability and its Applications, 61 (3):375–388, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Magdziarz:2010:NMM

[Mag10] M. Magdziarz. A note on Maruyama’s mixing theorem. Theory of Probability and its Applications, 54(2):322–324, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i2/p322_s1.

Makasu:2017:MEI

[Mak17] C. Makasu. Maximal exponential inequalities for certain diffusion processes. Theory of Probability and its Applications, 61 (1):159–167, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malyshev:2017:DDC

[Mal17] F. M. Malyshev. The distribution density of the commutant of random rotations of a threedimensional Euclidean space. Theory of Probability and its Applications, 61(2):277–294, ???? 2017. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mallein:2018:BRW

[Mal18a] B. Mallein. N-branching random walk with α-stable spine. Theory of Probability and its Applications, 62(2):295–318, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maller:2018:ERL

[Mal18b] R. A. Maller. Extensions of regularity for a Lévy process. Theory of Probability and its Applications, 62(4):575–603, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mamporia:2012:SDE

[Mam12] B. Mamporia. Stochastic differential equation for generalized random processes in a Banach space. Theory of Probability and its Applications, 56(4):602–620, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mamporia:2013:SDE

[Mam13] B. Mamporia. Stochastic differential equation for generalized random processes in a Banach space. Theory of Probability and its Applications, 56(4):602–620, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mamporia:2017:LSD

[Mam17] B. Mamporia. Linear stochastic differential equations in a Banach space. Theory of Probability and its Applications, 61 (2):295–308, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Manita:2015:NSL

[Man15] A. D. Manita. News of scientific life. Theory of Probability and its Applications, 59 (4):707–710, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mastikhin:2012:FPB

[Mas12] A. V. Mastikhin. Final probabilities for Becker epidemic Markov process. Theory of Probability and its Applications, 56(3):521–527, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mazurkiewicz:2012:BWS

[Maz12] G. Mazurkiewicz. On R^{+}- Weakly stable distribution. Theory of Probability and its Applications, 56(1):149–154, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v56/i1/p149_s1.

Mazur:2018:MFT

[Maz18] A. E. Mazur. Modeling and fitting of time series with heavy distribution tails and strong time dependence by Gaussian time series. Theory of Probability and its Applications, 63(1):151–154, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Myronyuk:2015:ILS

[MF15] M. V. Myronyuk and G. M. Feldman. Independent linear statistics on the cylinders. Theory of Probability and its Applications, 59 (2):260–278, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maslov:2013:UPT

[MM13] V. P. Maslov and T. V. Maslova. Unbounded probability theory and its applications. Theory of Probability and its Applications, 57 (3):444–467, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mohamed:2014:MCS

[Moh14] N. M. A. Mohamed. Monte Carlo sampling of Maxwell and Gaussian distributions using a single random number. Theory of Probability and its Applications, 58 (4):698–704, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Molchanov:2012:NSL

[Mol12] S. A. Molchanov. News of scientific life: On the 80th birthday of A. A. Yushkevich. Theory of Probability and its Applications, 56(1):176–177, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v56/i1/p176_s1.

Morozov:2012:DVF

[Mor12] I. S. Morozov. Differentiability of the value function in the robust utility maximization problem. Theory of Probability and its Applications, 56(2):327–337, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Metafune:2010:GPT

[MPR10] G. Metafune, D. Pallara, and A. Rhandi. Global properties of transition probabilities of singular diffusions. Theory of Probability and its Applications, 54(1):68–96, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Modarresi:2015:CPC

[MR15] N. Modarresi and S. Rezakhah.

Certain periodically correlated multicomponent locally stationary processes. Theory of Probability and its Applications, 59 (2):320–327, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Meschenmoser:2013:FCL

[MS13] D. Meschenmoser and A. Shashkin.

Functional Central Limit Theorem for the measures of level surfaces of the Gaussian random field. Theory of Probability and its Applications, 57(1):162–172, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Melkumova:2019:MLM

[MS19] L. E. Melkumova and S. Ya. Shatskikh. Maximum likelihood method in de Finetti’s Theorem. Theory of Probability and its Applications, 63(4):657–663, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Muminov:2011:LDM

[Mum11] M. S. Muminov. On limit distribution of maximal deviation of empirical distribution density and regression function. I. Theory of Probability and its Applications, 55(3):509–517, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i3/p509_s1.

Muminov:2012:LDM

[Mum12] M. S. Muminov. On limit distribution of maximal deviation of empirical distribution density and regression function. II. Theory of Probability and its Applications, 56(1):155–166, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v56/i1/p155_s1.

Muravlev:2011:PDB

[Mur11] A. A. Muravlev. On a property of the distribution of a Brownian motion with drift and its maximum. Theory of Probability and its Applications, 55(2):308–315, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i2/p308_s1.

Mikulevicius:2018:WES

[MZ18] R. Mikulevicius and Ch. Zhang.

Weak Euler scheme for Lévydriven stochastic differential equations. Theory of Probability and its Applications, 63(2):246–266, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:2010:NPA

[Nag10] S. V. Nagaev. A new proof of the absolute convergence of the Spitzer series. Theory of Probability and its Applications, 54 (1):151–154, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:2011:SCS

[Nag11] S. V. Nagaev. On sufficient conditions for subexponentiality. Theory of Probability and its Applications, 55(1):153–164, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i1/p153_s1.

Nagaev:2012:RTA

[Nag12] S. V. Nagaev. The renewal theorem in the absence of power moments. Theory of Probability and its Applications, 56(1):166–175, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v56/i1/p166_s1.

Nagaev:2015:LRT

[Nag15a] S. V. Nagaev. Local renewal theorems in the absence of an expectation. Theory of Probability and its Applications, 59 (3):388–414, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:2015:PIG

[Nag15b] S. V. Nagaev. Probabilistic inequalities for the Galton–Watson processes. Theory of Probability and its Applications, 59(4): 611–640, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Novikov:2017:PAT

[NAKL17] A. Novikov, S. Alexander, N. Kordzakhia, and T. Ling. Pricing of Asian-type and basket options via bounds. Theory of Probability and its Applications, 61 (1):94–106, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nazarov:2010:STG

[Naz10] A. I. Nazarov. On a set of transformations of Gaussian random functions. Theory of Probability and its Applications, 54(2):203–216, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i2/p203_s1.

Nadaraya:2013:ISM

[NBS13] E. Nadaraya, P. Babilua, and G. Sokhadze. On the integral square measure of deviation of a nonparametric estimator of the Bernoulli regression. Theory of Probability and its Applications, 57 (2):265–278, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:2012:BPB

[NC12] S. V. Nagaev and V. I. Chebotarev. On the bound of proximity of the binomial distribution to the normal one. Theory of Probability and its Applications, 56 (2):213–239, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nezakati:2011:LTS

[Nez11] A. Nezakati. Limit theorems for sequences of blockwise negatively associated random variables. Theory of Probability and its Applications, 55(2):315–321, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i2/p315_s1.

Nikitin:2013:NSL

[Nik13] Ya. Yu. Nikitin. News of scientific life. Theory of Probability and its Applications, 57(3):545, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Novikov:2013:PEA

[NK13] A. Novikov and N. Kordzakhia. Pitman estimators: an asymptotic variance revisited. Theory of Probability and its Applications, 57(3):521–529, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Novikov:2014:MPE

[NKL14] A. Novikov, N. Kordzakhia, and T. Ling. On moments of Pitman estimators: The case of fractional Brownian motion. Theory of Probability and its Applications, 58(4):601–614, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nguyen-Ngoc:2011:LLP

[NN11]  L. Nguyen-Ngoc.  Limiting

laws and penalization of certain Lévy processes by a function of their maximum. Theory of Probability and its Applications, 55(3):444–459, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v55/i3/p444_s1.

Novikov:2012:LMP

[NN12] An. A. Novikov and P. A. Novikov. Locally most powerful sequential tests of a simple hypothesis vs. one-sided alternatives for independent observations. Theory of Probability and its Applications, 56 (3):420–442, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Noarov:2010:SDP

[Noa10] A. I. Noarov. On some diffusion processes with stationary distributions. Theory of Probability and its Applications, 54(3):525–533, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i3/p525_s1.

Novikov:2011:LMP

[Nov11] P. A. Novikov. Locally most powerful sequential tests for discretetime Markov processes. Theory of Probability and its Applications, 55(2):322–325, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i2/p322_s1.

Novak:2014:AIH

[Nov14] S. Y. Novak. On the accuracy of inference on heavy-tailed distributions. Theory of Probability and its Applications, 58 (3):509–518, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nikitin:2013:ESD

[NP13] Ya. Yu. Nikitin and R. S. Pusev. Exact small deviation asymptotics for some Brownian functionals. Theory of Probability and its Applications, 57(1):60–81, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nazarov:2016:SBA

[NP16] A. I. Nazarov and Y. P. Petrova. The small ball asymptotics in Hilbert norm for the Kac–Kiefer– Wolfowitz processes. Theory of Probability and its Applications, 60 (3):460–480, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nie:2016:BDM

[NR16] T. Nie and M. Rutkowski. BSDEs driven by multidimensional martingales and their applications to markets with funding costs. Theory of Probability and its Applications, 60(4):604–630, ???? 2016. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nefedova:2013:NCR

[NS13] Yu. S. Nefedova and I. G. Shevtsova. On nonuniform convergence rate estimates in the Central Limit Theorem. Theory of Probability and its Applications, 57(1):28–59, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nadaraya:2010:SGF

[NSB10] E. Nadaraya, G. Sokhadze, and P. Babilua. On some goodnessof-fit tests based on estimates of kernel-type distribution densities. Theory of Probability and its Applications, 54(2):324–333, ???? 2010. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i2/p324_s1.

ORourke:2014:FME

[ORS14] S. O’Rourke, D. Renfrew, and A. Soshnikov. Fluctuations of matrix entries of regular functions of sample covariance random matrices. Theory of Probability and its Applications, 58 (4):615–639, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Osmolovskii:2010:EAE

[Osm10] I. Yu. Osmolovskii. On estimating approximation exactness for asymptotic expansions in polynomial cases. Theory of Probability and its Applications, 54 (1):154–160, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Palamarchuk:2018:AAB

[Pal18] E. S. Palamarchuk. Analysis of the asymptotic behavior of the solution to a linear stochastic differential equation with subexponentially stable matrix and its application to a control problem. Theory of Probability and its Applications, 62(4):522–533, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Palamarchuk:2019:UFA

[Pal19] E. S. Palamarchuk. On upper functions for anomalous diffusions governed by time-varying Ornstein–Uhlenbeck process. Theory of Probability and its Applications, 64(2):209–228, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Perelman:2012:TVI

[Per12] G. V. Perelman. Towards the validity of Itô’s formula for discontinuous functions. Theory of Probability and its Applications, 56 (3):443–456, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Privalsky:2012:MPC

[PF12] V. E. Privalsky and M. I. Fortus. Miscellanea: On possible causes of global warming. Theory of Probability and its Applications, 56(2):313–317, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pancheva:2014:ABC

[PG14] E. I. Pancheva and A. Gacovska. Asymptotic behavior of central order statistics under monotone normalization. Theory of Probability and its Applications, 58 (1):107–120, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:2018:BCF

[PGU18] Yu. V. Prokhorov, F. Götze, and V. V. Ulyanov. On bounds for characteristic functions of powers of asymptotically normal variables. Theory of Probability and its Applications, 62(1):98–116, ???? 2018. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Pinelis:2015:RTI

[Pin15] I. Pinelis. Rosenthal-type inequalities for martingales in 2-smooth Banach spaces. Theory of Probability and its Applications, 59 (4):699–706, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pinelis:2018:QMN

[Pin18] I. Pinelis. Quantifying minimal noncollinearity among random points. Theory of Probability and its Applications, 62 (4):604–616, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pinelis:2019:EBT

[Pin19a] I. Pinelis. Exact bounds on the truncated-tilted mean, with applications. Theory of Probability and its Applications, 63 (3):447–463, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pinelis:2019:OUB

[Pin19b] I. Pinelis. An optimal upper bound on the tail probability for sums of random variables. Theory of Probability and its Applications, 64 (3):474–480, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Piterbarg:2018:MEG

[Pit18] V. I. Piterbarg. Massive excursions of Gaussian isotropic fields. method of moments. Theory of Probability and its Applications, 63 (2):193–208, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pang:2010:RCT

[PL10] T.-X. Pang and Z.-Y. Lin. On the rates of the Chung-type law of logarithm. Theory of Probability and its Applications, 54 (4):703–717, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Platonova:2017:PRS

[Pla17] M. V. Platonova. Probabilistic representation of a solution of the Cauchy problem for evolution equations with Riemann– Liouville operators. Theory of Probability and its Applications, 61 (3):389–407, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Polunchenko:2019:NQS

[PMS19] A. S. Polunchenko, S. Martínez, and J. San Martín. A note on the quasi-stationary distribution of the Shiryaev martingale on the positive half-line. Theory of Probability and its Applications, 63 (3):464–478, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pavlov:2015:OST

[PN15] I. V. Pavlov and O. V. Nazarko. Optional sampling theorem for deformed submartingales. Theory of Probability and its Applications, 59(3):499–507, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Polunchenko:2018:ANM

[Pol18]  A. S. Polunchenko.  Asymp-

totic near-minimaxity of the randomized Shiryaev–Roberts–Pollak change-point detection procedure in continuous time. Theory of Probability and its Applications, 62 (4):617–631, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pozdnyakov:2013:BI

[Poz13] V. Pozdnyakov. On the 80th birthday of I. A. Ibragimov. Theory of Probability and its Applications, 57(3):367, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Peccati:2010:DSP

[PP10] G. Peccati and J.-R. Pycke.

Decompositions of stochastic processes based on irreducible group representations. Theory of Probability and its Applications, 54(2):217–245, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i2/p217_s1.

Percus:2017:JSR

[PP17] J. K. Percus and O. E. Percus. Joint statistics of random walk on Z^{1} and accumulation of visits. Theory of Probability and its Applications, 61 (3):499–505, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokaj:2011:LTM

[PR11] V. Prokaj and M. Rásonyi. Local and true martingales in discrete time. Theory of Probability and its Applications, 55(2):325–332, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v55/i2/p325_s1.

Platonova:2019:BRW

[PR19] M. V. Platonova and K. S. Ryadovkin. Branching random walks on Z^{d} with periodic branching sources. Theory of Probability and its Applications, 64 (2):229–248, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:2018:YPA

[Pro18] Yuri Prokhorov. Yuri Prokhorov about himself. Theory of Probability and its Applications, 62(1):2– 11, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Przystalski:2015:NLI

[Prz15] M. Przystalski. A note on a law of iterated logarithm for bounded Ndemimartingales. Theory of Probability and its Applications, 59 (1):163–167, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Przystalski:2016:LE

[Prz16] M. Przystalski. Letter to the Editors. Theory of Probability and its Applications, 60(3):528, ???? 2016. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Piterbarg:2010:NSLa

[PS10a] V. Piterbarg and A. N. Shiryaev.

News of scientific life. Theory of Probability and its Applications, 54(2):370–373, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i2/p370_s1.

Piterbarg:2010:NSLb

[PS10b] V. Piterbarg and A. N. Shiryaev.

News of scientific life. Theory of Probability and its Applications, 54(3):550–551, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i3/p550_s1.

Presman:2010:OSR

[PS10c] E. L. Presman and I. M. Sonin.

On optimal stopping of random sequences modulated by Markov chain. Theory of Probability and its Applications, 54(3):534–542, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i3/p534_s1.

Pirogov:2014:NSL

[PS14] S. A. Pirogov and R. Sverchkov. News of scientific life. Theory of Probability and its Applications, 58 (4):705–707, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:2014:BBV

[PSM14] Yu. V. Prokhorov, A. N. Shiryaev, and A. D. Manita. On the 100th birthday of B. V. Gnedenko. Theory of Probability and its Applications, 58(1):1–3, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pavlov:2017:EMM

[PST17] I. V. Pavlov, V. V. Shamrayeva, and I. V. Tsvetkova. On the existence of martingale measures satisfying the weakened condition of noncoincidence of barycenters in the case of countable probability space. Theory of Probability and its Applications, 61 (1):167–175, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pusev:2011:SDA

[Pus11] R. S. Pusev. Small deviations asymptotics for Matérn processes and fields under weighted quadratic norm. Theory of Probability and its Applications, 55(1):164–172, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i1/p164_s1.

Ragimov:2013:ILT

[RA13] F. Ragimov and F. Azizov. Integral limit theorems for nonlinear boundary crossing time for Markov chain. Theory of Probability and its Applications, 57 (1):172–178, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Radchenko:2010:SFS

[Rad10] V. N. Radchenko. Sample functions of stochastic measures and Besov spaces. Theory of Probability and its Applications, 54 (1):160–168, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Radchenko:2015:EED

[Rad15] V. M. Radchenko. Evolution equations driven by general stochastic measures in Hilbert space. Theory of Probability and its Applications, 59(2):328–339, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Radchenko:2018:FSE

[Rad18] V. M. Radchenko. Fourier series expansion of stochastic measures. Theory of Probability and its Applications, 63(2):318–326, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ragimov:2010:ABL

[Rag10] F. G. Ragimov. On the asymptotic behavior of local probabilities of a multidimensional random walk crossing nonlinear boundaries. Theory of Probability and its Applications, 54(2):333–339, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i2/p333_s1.

Raimova:2013:PRS

[Rai13] G. M. Raimova. Probabilistic representation of the solution of the initial-boundary value problem for the system of parabolic equations. Theory of Probability and its Applications, 57(4): 688–697, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rao:2018:CPD

[Rao18] B. L. S. Prakasa Rao. Characterizations of probability distributions through Q-independence. Theory of Probability and its Applications, 62(2):335–338, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ruiz-Medina:2014:CLR

[RMC14] M. D. Ruiz-Medina and R. M. Crujeiras. A central limit result in the wavelet domain for minimum contrast estimation of fractal random fields. Theory of Probability and its Applications, 58 (3):458–486, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rodionov:2018:DTT

[Rod18] I. V. Rodionov. A discrimination test for tails of Weibull-type distributions. Theory of Probability and its Applications, 63 (2):327–335, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rodionov:2019:DCH

[Rod19] I. V. Rodionov. Discrimination of close hypotheses about the distribution tails using higher order statistics. Theory of Probability and its Applications, 63 (3):364–380, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rokhlin:2010:LBM

[Rok10] D. B. Rokhlin. Lower bounds of martingale measure densities in the Dalang–Morton– Willinger theorem. Theory of Probability and its Applications, 54(3):447–465, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i3/p447_s1.

Rokhlin:2012:RRP

[Rok12] D. B. Rokhlin. Recurrence relations for price bounds of contingent claims in discrete time market models. Theory of Probability and its Applications, 56(1):72–95, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v56/i1/p72_s1.

Rokhlin:2019:LSS

[Rok19] D. B. Rokhlin. Q-learning in a stochastic Stackelberg game between an uninformed leader and a naive follower. Theory of Probability and its Applications, 64(1):41– 58, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovsky:2010:SDP

[Roz10] L. V. Rozovsky. Small deviation probabilities for maximum of partial sums. Theory of Probability and its Applications, 54 (4):717–724, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovsky:2013:CCG

[Roz13] L. V. Rozovsky. On the convergence conditions of the generalized Spitzer series. Theory of Probability and its Applications, 57 (4):697–698, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovsky:2014:SDS

[Roz14] L. V. Rozovsky. Small deviations of series of independent nonnegative random variables with smooth weights. Theory of Probability and its Applications, 58 (1):121–137, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovsky:2015:SDP

[Roz15] L. V. Rozovsky. Superlarge deviation probabilities for sums of independent random variables with exponentially decreasing tails. II. Theory of Probability and its Applications, 59(1):168–177, ???? 2015. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovsky:2016:SDP

[Roz16] L. V. Rozovsky. Small deviations of probabilities for weighted sum of independent positive random variables with a common distribution that decreases at zero not faster than a power. Theory of Probability and its Applications, 60(1):142–150, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovsky:2018:SDPb

[Roz18a] L. V. Rozovsky. Small deviation probabilities for a weighted sum of independent positive random variables with common distribution function that can decrease at zero fast enough. Theory of Probability and its Applications, 63 (1):155–163, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovsky:2018:SDPa

[Roz18b] L. V. Rozovsky. Small deviation probabilities for weighted sum of independent random variables with a common distribution having a power decrease at zero, under minimal moment assumptions. Theory of Probability and its Applications, 62(3): 491–495, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovsky:2019:ILC

[Roz19a] L. V. Rozovsky. Integro-local CLT for sums of independent nonlattice random vectors. Theory of Probability and its Applications, 64(1):27–40, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovsky:2019:EAS

[Roz19b] L. V. Rozovsky. On the exact asymptotics of small deviations of L_{2}-norm for some Gaussian random fields. Theory of Probability and its Applications, 63 (3):381–392, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rasonyi:2016:OIU

[RRV16] M. Rásonyi and J. G. Rodríguez-Villarreal. Optimal investment under behavioral criteria in incomplete diffusion market models. Theory of Probability and its Applications, 60(4):631–646, ???? 2016. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rukhin:2010:JDP

[Ruk10] A. L. Rukhin. Joint distribution of pattern frequencies and multivariate Pòlya–Aeppli law. Theory of Probability and its Applications, 54(2):246–260, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v54/i2/p246_s1.

Sakhanenko:2010:LBA

[Sak10] L. Sakhanenko. Lower bounds for accuracy of estimation in diffusion tensor imaging. Theory of Probability and its Applications, 54 (1):168–177, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sakhanenko:2011:GOC

[Sak11] L. Sakhanenko. A global optimal convergence rate in a model for the diffusion tensor imaging. Theory of Probability and its Applications, 55(1):77–90, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i1/p77_s1.

Sakhanenko:2015:ASW

[Sak15] L. Sakhanenko. Asymptotics of suprema of weighted Gaussian fields with applications to kernel density estimators. Theory of Probability and its Applications, 59(3):415–451, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sakhanenko:2019:TMD

[Sak19] L. A. Sakhanenko. Testing a multivariate distribution for generalized skew ellipticity. Theory of Probability and its Applications, 64 (2):290–303, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sauri:2019:PDB

[Sau19] O. Sauri. Pathwise decompositions of Brownian semistationary processes. Theory of Probability and its Applications, 64 (1):78–102, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Savinov:2015:LTM

[Sav15] E. A. Savinov. Limit theorem for the maximum of random variables connected by IT-copulas of Student’s t-distribution. Theory of Probability and its Applications, 59(3):508–516, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Savelov:2017:ECS

[Sav17] M. P. Savelov. Extremal characteristics of statistical criteria with given total variation distances between hypotheses. Theory of Probability and its Applications, 61 (3):442–461, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Savelov:2019:GIS

[Sav19] M. P. Savelov. Gittins index for simple family of Markov bandit processes with switching cost and no discounting. Theory of Probability and its Applications, 64 (3):355–364, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Simas:2012:ATP

[SCN12] A. B. Simas, G. M. Cordeiro, and S. Nadarajah. Asymptotic tail properties of the distributions in the class of dispersion models. Theory of Probability and its Applications, 56(4):703–710, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Simas:2013:ATP

[SCN13] A. B. Simas, G. M. Cordeiro, and S. Nadarajah. Asymptotic tail properties of the distributions in the class of dispersion models. Theory of Probability and its Applications, 56(4):703–710, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Semenov:2019:TCR

[Sem19] A. S. Semenov. Two-colorings of a random hypergraph. Theory of Probability and its Applications, 64(1):59–77, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Senatov:2010:AEL

[Sen10a] V. V. Senatov. On approximation exactness in the law of large numbers. Theory of Probability and its Applications, 54(2):261–271, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i2/p261_s1.

Senatov:2010:MAC

[Sen10b] V. V. Senatov. On multivariate analogues of Chebyshev–Hermite polynomials. Theory of Probability and its Applications, 54 (4):659–676, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Senatov:2011:QNE

[Sen11] V. V. Senatov. On quasinonuniform estimates of the approximation exactness in the Central Limit Theorem. Theory of Probability and its Applications, 55(3):460–472, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v55/i3/p460_s1.

Senatov:2012:OAE

[Sen12] V. V. Senatov. On one asymptotic expansion in the Central Limit Theorem. Theory of Probability and its Applications, 56 (2):337–343, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Senatov:2013:NEA

[Sen13] V. V. Senatov. Nonimprovable estimates for asymptotic expansions in the Central Limit Theorem. Theory of Probability and its Applications, 57(4):633–648, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Senatov:2015:NEA

[Sen15] V. V. Senatov. On nonuniform estimates of approximation exactness in the central limit theorem. Theory of Probability and its Applications, 59(2):279–310, ???? 2015. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Senatov:2016:CHP

[Sen16] V. V. Senatov. On construction of high precision asymptotic expansions. Theory of Probability and its Applications, 60 (2):295–311, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Serdyukova:2010:DDC

[Ser10] Nora A. Serdyukova. Dependence on the dimension for complexity of approximation of random fields. Theory of Probability and its Applications, 54(2):272–284, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i2/p272_s1.

Serov:2011:LDH

[Ser11] A. A. Serov. Limit distribution of the Hamming distance from the random Boolean function to the set of affine functions. Theory of Probability and its Applications, 55(4):717–722, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i4/p717_s1.

Shepp:2011:ENR

[SF11] L. Shepp and K. Farahmand. Expected number of real zeros of a random polynomial with independent identically distributed symmetric long-tailed coefficients. Theory of Probability and its Applications, 55(1):173–181, ????

2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v55/i1/p173_s1.

Shaposhnikov:2012:SUS

[Sha12a] A. V. Shaposhnikov. On a strong uniqueness of solutions of degenerate stochastic differential equations. Theory of Probability and its Applications, 56(2):240–251, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shaposhnikov:2012:UPS

[Sha12b] S. V. Shaposhnikov. On uniqueness of a probability solution to the Cauchy problem for the Fokker–Planck–Kolmogorov equation. Theory of Probability and its Applications, 56(1):96–115, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v56/i1/p96_s1.

Shaposhnikov:2012:RQP

[Sha12c] S. V. Shaposhnikov. Regularity and qualitative properties of solutions of parabolic equations for measures. Theory of Probability and its Applications, 56 (2):252–279, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shashkin:2015:ANL

[Sha15] A. Shashkin. Asymptotic normality of local averaging estimates for weakly dependent random fields. Theory of Probability and its Applications, 59 (3):516–526, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shashkin:2016:FLT

[Sha16] A. Shashkin. A functional limit theorem for the integrals over level sets of a Gaussian random field. Theory of Probability and its Applications, 60(1):150–161, ???? 2016. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shevtsova:2011:AEC

[She11] I. G. Shevtsova. On the asymptotically exact constants in the Berry–Esseen-katz inequality. Theory of Probability and its Applications, 55(2):225–252, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i2/p225_s1.

Shevtsova:2013:MTE

[She13] I. G. Shevtsova. Moment-type estimates with an improved structure for the accuracy of the normal approximation to distributions of sums of independent symmetric random variables. Theory of Probability and its Applications, 57 (3):468–496, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shevtsova:2014:NSL

[She14a] I. G. Shevtsova. News of scientific life. Theory of Probability and its Applications, 58 (2):343–344, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shevtsova:2014:ANA

[She14b] I. G. Shevtsova. On the accuracy of the normal approximation to compound Poisson distributions. Theory of Probability and its Applications, 58(1): 138–158, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shevtsova:2018:CRE

[She18a] I. G. Shevtsova. Convergence rate estimates in the global CLT for compound mixed Poisson distributions. Theory of Probability and its Applications, 63(1):72–93, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shevtsova:2018:MIA

[She18b] I. G. Shevtsova. A moment inequality with application to convergence rate estimates in the global CLT for Poisson-binomial random sums. Theory of Probability and its Applications, 62 (2):278–294, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2011:NSL

[Shi11] A. N. Shiryaev. News of scientific life. Theory of Probability and its Applications, 55(2):368–370, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i2/p368_s1.

Shiryaev:2015:NSLb

[Shi15a] A. Shiryaev. News of scientific life. Theory of Probability and its Applications, 59(2):355–356, ???? 2015. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2015:NSLa

[Shi15b] A. N. Shiryaev. News of scientific life. Theory of Probability and its Applications, 59 (1):178–179, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2017:MOC

[Shi17a] A. N. Shiryaev. On the minimax optimality of the CUSUM statistic in disorder problems for Brownian motion. Theory of Probability and its Applications, 61 (4):719–726, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2017:OR

[Shi17b] A. N. Shiryaev. To our readers!

Theory of Probability and its Applications, 61(4):539, ???? 2017. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2018:NSL

[Shi18a] A. N. Shiryaev. News of scientific life-information on the general seminar of the department of probability (faculty of mathematics and mechanics, Moscow state university). Theory of Probability and its Applications, 62 (4):675–680, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2018:BAI

[Shi18b] A. N. Shiryaev. On the 85th birthday of Academician I. A. Ibragimov. Theory of Probability and its Applications, 62(4):505, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shin:2019:SDH

[Shi19a] J. Shin. Semimartingale decomposition and heat kernel estimates of reflected stable-like processes with variable order. Theory of Probability and its Applications, 64 (3):421–443, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2019:MRS

[Shi19b] A. N. Shiryaev. In memory of R. Sh. Liptser (20.03.1936– 02.01.2019). Theory of Probability and its Applications, 64(2):334, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shklyaev:2011:LDS

[Shk11a] A. V. Shklyaev. Large deviations of Shepp statistics. Theory of Probability and its Applications, 55(4):722–729, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i4/p722_s1.

Shklyaev:2011:LTR

[Shk11b] A. V. Shklyaev. Limit theorems for random walk under the assumption of maxima large deviation. Theory of Probability and its Applications, 55(3):517–525, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i3/p517_s1.

Shur:2011:UIS

[Shu11] M. G. Shur. Uniform integrability for strong ratio limit theorems. II. Theory of Probability and its Applications, 55(3):473–484, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i3/p473_s1.

Shur:2012:LE

[Shu12] M. G. Shur. Letters to the Editors. Theory of Probability and its Applications, 56(1):178–179, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v56/i1/p178_s1.

Shur:2013:UIS

[Shu13] M. G. Shur. Uniform integrability for strong ratio limit theorems. III. Theory of Probability and its Applications, 57 (4):649–662, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:2014:TTC

[Shu14] M. G. Shur. Two theorems on convergence parameter of an irreducible Markov chain. Theory of Probability and its Applications, 58(1):159–164, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shur:2017:ERR

[Shu17] M. G. Shur. Exponentials and R-recurrent random walks on groups. Theory of Probability and its Applications, 61(3):505–513, ???? 2017. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shvedov:2015:MVD

[Shv15] A. S. Shvedov. Matrix-variate t-distribution with vector of degrees of freedom. Theory of Probability and its Applications, 59 (3):526–531, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sidorov:2010:MCS

[Sid10] D. I. Sidorov. On mixing conditions for sequences of moving averages. Theory of Probability and its Applications, 54(2):339–347, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v54/i2/p339_s1.

Sinelnikov:2012:OSB

[Sin12] S. S. Sinelnikov. On the optimal stopping of a Brownian motion with a negative drift. Theory of Probability and its Applications, 56(2):343–350, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sinn:2011:CZC

[SK11] M. Sinn and K. Keller. Covariances of zero crossings in Gaussian processes. Theory of Probability and its Applications, 55(3):485–504, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i3/p485_s1.

Stoyanov:2013:HCM

[SL13] J. Stoyanov and G. D. Lin.

Hardy’s condition in the moment problem for probability distributions. Theory of Probability and its Applications, 57(4):699–708, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sokolov:2014:RCP

[Sok14] N. V. Sokolov. Refinements of classical probability estimates. Theory of Probability and its Applications, 58(2):331–333, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Song:2016:RTD

[Son16] Sh. Song. Random time with differentiable conditional distribution function. Theory of Probability and its Applications, 60 (4):647–669, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Stepanova:2018:GFT

[SP18] N. Stepanova and T. Pavlenko. Goodness-of-fit tests based on supfunctionals of weighted empirical processes. Theory of Probability and its Applications, 63 (2):292–317, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2019:ATG

[SP19] A. N. Shiryaev and I. V. Pavlov. Abstracts of talks given at the 3rd International Conference on Stochastic Methods. Theory of Probability and its Applications, 64 (1):124–169, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Spiryaev:2012:SPK

[Spi12] M. A. Spiryaev. On some properties of Kagi and Renko trading strategies for Brownian motion. Theory of Probability and its Applications, 56(2):280–298, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2017:AIC

[SPTS17] A. N. Shiryaev, I. V. Pavlov, T. B. Tolozova, and V. V. Shamraeva. Abstracts of the International Conference on Stochastic Methods. Theory of Probability and its Applications, 61 (3):521–538, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Senatov:2013:NFA

[SS13] V. V. Senatov and V. N. Sobolev.

New forms of asymptotic expansions in the Central Limit Theorem. Theory of Probability and its Applications, 57(1):82–96, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Suvorikova:2017:MCP

[SS17] A. Suvorikova and V. Spokoiny.

Multiscale change point detection. Theory of Probability and its Applications, 61(4):665–691, ???? 2017. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Startsev:2015:OAE

[Sta15] A. N. Startsev. On one approach to estimation of parameters of a two-dimensional process of linear diffusion in nonstationary case. Theory of Probability and its Applications, 59 (2):339–343, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2011:RJS

[SU11] A. N. Shiryaev and V. V.

Ulyanov. Russian–Japanese symposium on “A stochastic analysis of advanced statistical models”. Theory of Probability and its Applications, 55(3):536–557, ???? 2011. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v55/i3/p536_s1.

Su:2013:NSL

[Su13] Zhonggen Su. News of scientific life. Theory of Probability and its Applications, 57(1): 179–180, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Su:2015:NCR

[Su15] Z. Su. Normal convergence for random partitions with multiplicative measures. Theory of Probability and its Applications, 59(1):40–69, ???? 2015. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sukhanova:2010:MC

[Suk10] E. M. Sukhanova. Matrix correlation. Theory of Probability and its Applications, 54(2):347–355, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v54/i2/p347_s1.

Salminen:2018:EFP

[SV18] P. Salminen and L. Vostrikova. On exponential functionals of processes with independent increments. Theory of Probability and its Applications, 63(2):267–291, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shneer:2011:UAH

[SW11] S. Shneer and V. Wachtel. A unified approach to the heavytraffic analysis of the maximum of random walks. Theory of Probability and its Applications, 55(2):332–341, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i2/p332_s1.

Syulyukin:2010:ABC

[Syu10a] A. V. Syulyukin. On asymptotic Bergström-Chebyshev expansions. Theory of Probability and its Applications, 54(1): 177–187, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Syulyukin:2010:QNE

[Syu10b] A. V. Syulyukin. On quasinonuniform estimates of residuals for the Bergström–Chebyshev expansion. Theory of Probability and its Applications, 54(2):355–363, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i2/p355_s1.

Syulyukin:2014:AEC

[Syu14] A. V. Syulyukin. On asymptotic expansions for convolution of distributions which belong to stable laws. Theory of Probability and its Applications, 58 (3):518–524, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Szablowski:2012:WOU

[Sza12] P. J. Szabłowski. q-wiener and (α, q)-Ornstein–Uhlenbeck processes. A generalization of known processes. Theory of Probability and its Applications, 56(4): 634–659, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Szablowski:2013:WOU

[Sza13] P. J. Szabłowski. q-Wiener and (α, q)-Ornstein–Uhlenbeck processes. A generalization of known processes. Theory of Probability and its Applications, 56(4): 634–659, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Satin:2013:RCT

[SZK13] Ya. A. Satin, A. I. Zeifman, and A. V. Korotysheva. On the rate of convergence and truncations for a class of Markovian queueing systems. Theory of Probability and its Applications, 57 (3):529–539, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tella:2016:PRP

[TE16] P. Di Tella and H.-J. Engelbert. The predictable representation property of compensatedcovariation stable families of martingales. Theory of Probability and its Applications, 60(1):19–44, ???? 2016. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Timashev:2017:RPP

[Tim17] A. N. Timashev. Random permutations with prime lengths of cycles. Theory of Probability and its Applications, 61(2):309–320, ???? 2017. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Timashev:2018:LTP

[Tim18] A. N. Timashev. Limit theorems for power-series distributions with finite radius of convergence. Theory of Probability and its Applications, 63(1):45–56, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Timashev:2019:RMC

[Tim19] A. N. Timashev. Random mappings with component sizes from a given set. Theory of Probability and its Applications, 64 (3):481–489, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tolstikhin:2017:CIS

[Tol17] I. O. Tolstikhin. Concentration inequalities for samples without replacement. Theory of Probability and its Applications, 61 (3):462–481, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Topchii:2010:TEG

[Top10] V. A. Topchii. Time to extinction of Galton–Watson processes censored at high level. Theory of Probability and its Applications, 54(2):285–299, ????

2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i2/p285_s1.

Tartakovsky:2012:TOA

[TPP12] A. G. Tartakovsky, M. Pollak, and A. S. Polunchenko. Third-order asymptotic optimality of the generalized Shiryaev–Roberts changepoint detection procedures. Theory of Probability and its Applications, 56(3):457–484, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tyurin:2010:ULA

[Tyu10] I. S. Tyurin. Uniform laws of averages for sequences of dependent random elements. Theory of Probability and its Applications, 54(4):725–732, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tyurin:2011:CRL

[Tyu11a] I. S. Tyurin. On the convergence rate in Lyapunov’s theorem. Theory of Probability and its Applications, 55(2):253–270, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i2/p253_s1.

Tyurin:2011:MSA

[Tyu11b] Yu. N. Tyurin. Multivariate statistical analysis: The geometric theory. Theory of Probability and its Applications, 55(1):91–109, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i1/p91_s1.

Tyurin:2012:RRL

[Tyu12] I. S. Tyurin. A refinement of the remainder in the Lyapunov theorem. Theory of Probability and its Applications, 56(4): 693–696, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tyurin:2013:RRL

[Tyu13] I. S. Tyurin. A refinement of the remainder in the Lyapunov theorem. Theory of Probability and its Applications, 56(4): 693–696, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ulyanov:2016:PPR

[Uly16] V. V. Ulyanov. On properties of polynomials in random elements. Theory of Probability and its Applications, 60(2):325–336, ???? 2016. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vakeroudis:2012:HTW

[Vak12] S. Vakeroudis. On hitting times of the winding processes of planar Brownian motion and of Ornstein–Uhlenbeck processes via Bougerol’s identity. Theory of Probability and its Applications, 56 (3):485–507, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Varron:2015:NWC

[Var15] D. Varron. A note on weak convergence, large deviations, and the bounded approximation property. Theory of Probability and its Applications, 59(1):70–86, ???? 2015. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Varron:2019:ECM

[Var19] D. Varron. An extended continuous mapping theorem for outer almost sure weak convergence. Theory of Probability and its Applications, 64(2):304–323, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:2011:BYV

[Vat11a] V. Vatutin. On the 80th birthday of Yu. V. Prokhorov. Theory of Probability and its Applications, 55(1):1, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v55/i1/p1_s1.

Vatutin:2011:NSL

[Vat11b] V. A. Vatutin. News of scientific life. Theory of Probability and its Applications, 55(4):730–734, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v55/i4/p730_s1.

Vatutin:2011:PSM

[Vat11c] V. A. Vatutin. Polling systems and multitype branching processes in random environment with final product. Theory of Probability and its Applications, 55(4):631–660, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i4/p631_s1.

Vatutin:2015:SDR

[Vat15] V. A. Vatutin. The structure of the decomposable reduced branching processes. I. Finite-dimensional distributions. Theory of Probability and its Applications, 59 (4):641–662, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:2016:SDR

[Vat16] V. A. Vatutin. The structure of decomposable reduced branching processes. II. Functional limit theorems. Theory of Probability and its Applications, 60(1): 103–119, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vakhania:2010:QGR

[VC10] N. N. Vakhania and G. Z.

Chelidze. Quaternion Gaussian random variables. Theory of Probability and its Applications, 54(2):363–369, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i2/p363_s1.

Vatutin:2017:HMF

[VD17] V. A. Vatutin and E. E. Dyakonova. How many families survive for a long time? Theory of Probability and its Applications, 61(4):692–711, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:2018:EMB

[VD18a] V. A. Vatutin and E. E. Dyakonova. Erratum: Multitype Branching Processes in Random Environment: Probability of Survival for the Critical Case. Theory of Probability and its Applications, 63(2):338, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). See [VD18b].

Vatutin:2018:MBP

[VD18b] V. A. Vatutin and E. E. Dyakonova. Multitype branching processes in random environment: Probability of survival for the critical case. Theory of Probability and its Applications, 62(4):506–521, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). See [VD18a].

vanDoorn:2010:BAR

[vDZP10] E. A. van Doorn, A. I. Zeifman, and T. L. Panfilova. Bounds and asymptotics for the rate of convergence of birth-death processes. Theory of Probability and its Applications, 54(1):97–113, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Viskov:2016:FAY

[VK16] O. V. Viskov and V. I. Khokhlov.

Four areas of Yu. V. Prokhorov’s studies and their perspectives. Theory of Probability and its Applications, 60(2):336–342, ???? 2016. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vellaisamy:2018:FDE

[VK18] P. Vellaisamy and K. K. Kataria.

The I-function distribution and its extensions. Theory of Probability and its Applications, 63 (2):227–245, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:2013:CBP

[VL13] V. A. Vatutin and Q. Liu. Critical branching process with two types of particles evolving in asynchronous random environments. Theory of Probability and its Applications, 57(2):279–305, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volodko:2011:LTC

[Vol11] N. V. Volodko. Limit theorems for canonical von Mises and Ustatistics of m-dependent observations. Theory of Probability and its Applications, 55(2):271–290, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i2/p271_s1.

Vostrikova:2017:EUM

[Vos17] L. Vostrikova. Expected utility maximization for exponential Lévy models with option and information processes. Theory of Probability and its Applications, 61 (1):107–128, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Volodin:2010:IAB

[VP10] I. N. Volodin and M. N. Ponfilyenok. Improving approximation of the beta-distribution for small parameter values. Theory of Probability and its Applications, 54(3):543–549, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

                                        epubs.siam.org/tvp/resource/
                                        1/tprbau/v54/i3/p543_s1.

Vatutin:2014:KRT

[VT14] V. Vatutin and V. Topchii. A key renewal theorem for heavy tail distributions with β ∈ (0, 0.5]. Theory of Probability and its Applications, 58(2):333–342, ???? 2014. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vellaisamy:2013:NBA

[VUC13] P. Vellaisamy, N. S. Upadhye, and V. Cekanavicius. On negative binomial approximation. Theory of Probability and its Applications, 57(1):97–109, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:2010:SEC

[VW10] V. A. Vatutin and V. Wachtel. Sudden extinction of a critical branching process in a random environment. Theory of Probability and its Applications, 54(3):466–484, ???? 2010. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i3/p466_s1.

Vatutin:2016:MBS

[VZ16] V. A. Vatutin and A. M. Zubkov.

In memory of B. A. Sevast’yanov (29.09.1923–30.08.2013). Theory of Probability and its Applications, 60(1):162–171, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Wachtel:2016:EAM

[WD16] V. I. Wachtel and D. E. Denisov.

An exact asymptotics for the moment of crossing a curved boundary by an asymptotically stable random walk. Theory of Probability and its Applications, 60 (3):481–500, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Wakaki:2018:CEB

[WF18] H. Wakaki and Ya. Fujikoshi.

Computable error bounds for highdimensional approximations of an LR statistic for additional information in canonical correlation analysis. Theory of Probability and its Applications, 62 (1):157–172, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Wang:2014:IBS

[WH14] H. Wang and W. Hong. Intrinsic branching structure within random walk on Z. Theory of Probability and its Applications, 58 (4):640–659, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Wang:2019:BEB

[WH19] X. J. Wang and S. H. Hu. The Berry–Esseen bound for ρ-mixing random variables and its applications in nonparametric regression model. Theory of Probability and its Applications, 63 (3):479–499, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Wang:2018:MIN

[WHV18] X. J. Wang, S. H. Hu, and A. I. Volodin. Moment inequalities for m-NOD random variables and their applications. Theory of Probability and its Applications, 62(3):471–490, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Withers:2011:MCV

[WN11] C. S. Withers and S. Nadarajah. Maximum of continuous versions of Poisson and negative binomial-type distributions. Theory of Probability and its Applications, 55(3):525–528, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i3/p525_s1.

Wu:2015:SLT

[Wu15] Y. F. Wu. Some limit theorems for arrays of rowwise pairwise negatively quadratic dependent random variables. Theory of Probability and its Applications, 59(2):344–354, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yakymiv:2010:LTM

[Yak10] A. L. Yakymiv. Limit theorem for the middle members of ordered cycle lengths in random Apermutations. Theory of Probability and its Applications, 54 (1):114–128, ???? 2010. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yakymiv:2011:AIN

[Yak11] A. L. Yakymiv. Asymptotics at infinity of negative binomial infinitely divisible distributions. Theory of Probability and its Applications, 55(2):342–351, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i2/p342_s1.

Yakymiv:2015:NCR

[Yak15] A. L. Yakymiv. On a number of components in a random A-mapping. Theory of Probability and its Applications, 59 (1):114–127, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yakymiv:2016:TTG

[Yak16] A. L. Yakymiv. Tauberian theorem for generating functions of multiple series. Theory of Probability and its Applications, 60 (2):343–347, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yakymiv:2018:ORP

[Yak18] A. L. Yakymiv. On the order of random permutation with cycle weights. Theory of Probability and its Applications, 63 (2):209–226, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yakymiv:2019:ATR

[Yak19] A. L. Yakymiv. Abelian theorem for the regularly varying measure and its density in orthant. Theory of Probability and its Applications, 64(3):385–400, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yamazato:2012:ACC

[Yam12] M. Yamazato. Absolute continuity conditions for multivariate infinitely divisible distributions and their applications. Theory of Probability and its Applications, 56 (2):299–312, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yanushkevichiene:2012:BLT

[Yan12] O. Yanushkevichiene. On bounds in limit theorems for some Ustatistics. Theory of Probability and its Applications, 56(4):

660–673, ????  2012.  CO-

DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yanushkevichiene:2013:BLT

[Yan13] O. Yanushkevichiene. On bounds in limit theorems for some Ustatistics. Theory of Probability and its Applications, 56(4): 660–673, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yarovaya:2011:CEG

[Yar11] E. B. Yarovaya. Criteria of exponential growth for the numbers of particles in models of branching random walks. Theory of Probability and its Applications, 55(4):661–682, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i4/p661_s1.

Yarovaya:2018:SAS

[Yar18] E. B. Yarovaya. Spectral asymptotics of a supercritical branching random walk. Theory of Probability and its Applications, 62 (3):413–431, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yaskov:2010:BCR

[Yas10] P. A. Yaskov. A bound of convergence rate in weak law of large numbers for epidemic process. Theory of Probability and its Applications, 54(3):485–498, ???? 2010. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v54/i3/p485_s1.

Yaskov:2011:SCM

[Yas11] P. A. Yaskov. Strong convergence for multiple sums of nonorthogonal random variables. Theory of Probability and its Applications, 55(2):351–355, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i2/p351_s1.

Yaskov:2015:RMW

[Yas15] P. Yas’kov. Russian mathematician wins the 2014 Abel Prize. Theory of Probability and its Applications, 59(1):1–2, ???? 2015. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yaskov:2016:MBD

[Yas16] P. Yas’kov. In memory of E. B. Dynkin (11.05.1924–14.11.2014). Theory of Probability and its Applications, 60(1):172, ???? 2016. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yaskov:2018:SSC

[Yas18] P. A. Yaskov. On the spectrum of sample covariance matrices for time series. Theory of Probability and its Applications, 62 (3):432–443, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yengibaryan:2011:RWM

[YB11] N. B. Yengibaryan and A. G.

Barseghyan. Random walks and mixtures of gamma distributions. Theory of Probability and its Applications, 55(3):528–535, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i3/p528_s1.

Yilmaz:2015:SJD

[YMW15] Y. Yilmaz, G. V. Moustakides, and X. Wang. Sequential joint detection and estimation. Theory of Probability and its Applications, 59(3):452–465, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yongfeng:2012:CPM

[Yon12] Wu Yongfeng. Convergence properties of the maximal partial sums for arrays of rowwise NA random variables. Theory of Probability and its Applications, 56 (3):527–535, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yam:2013:UBB

[YYZ13] S. C. P. Yam, S. P. Yung, and W. Zhou. A unified “Bang– Bang” Principle with respect to R-invariant performance benchmarks. Theory of Probability and its Applications, 57(2):357–366, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zacharovas:2019:EDB

[Zac19] V. Zacharovas. The estimate of χ^{2}-distance between binomial and generalized binomial distributions. Theory of Probability and its Applications, 64(3):444–455, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zaikin:2015:DSN

[Zai15] A. A. Zaikin. Defect of the size of nonrandomized test and randomization effect on decreasing of the necessary sample size in testing the Bernoulli success probability. Theory of Probability and its Applications, 59 (3):466–480, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zaikin:2019:EAU

[Zai19] A. A. Zaikin. Estimates with asymptotically uniformly minimal d-risk. Theory of Probability and its Applications, 63(3):500–505, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zani:2013:SPL

[Zan13] M. Zani. Sample path large deviations for squares of stationary Gaussian processes. Theory of Probability and its Applications, 57 (2):347–357, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zineb:2017:AAV

[ZG17] T. Ben Zineb and E. Gobet. Analytical approximation of variable annuities for small volatility and small withdrawal. Theory of Probability and its Applications, 61 (1):40–56, ???? 2017. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhang:2012:BCF

[Zha12] Z. Zhang. Bounds for characteristic functions and Laplace transforms of probability distributions. Theory of Probability and its Applications, 56(2):350–358, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhdanov:2016:PHE

[Zhd16] A. I. Zhdanov. On probability of high extremes for product of two Gaussian stationary processes. Theory of Probability and its Applications, 60(3): 520–527, ???? 2016. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhukovskii:2011:WZO

[Zhu11] M. E. Zhukovskii. The weak zeroone law for the random distance graphs. Theory of Probability and its Applications, 55(2):356–360, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i2/p356_s1.

Zhukovskii:2014:LLN

[Zhu14] M. E. Zhukovskii. The law of large numbers for the number of active particles in an epidemic model. Theory of Probability and its Applications, 58 (2):297–313, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zubkov:2018:LTP

[ZK18] A. M. Zubkov and K. A. Kolesnikova. Limit theorem for the process of additive replacements. Theory of Probability and its Applications, 62(2):319–327, ???? 2018. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zubkov:2012:GBL

[ZKK12] A. M. Zubkov, N. A. Kharitonova, and E. V. Khil. Gaps between local maxima in sequences of random variables. Theory of Probability and its Applications, 56 (4):590–601, ???? 2012. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zubkov:2013:GBL

[ZKK13] A. M. Zubkov, N. A. Kharitonova, and E. V. Khil. Gaps between local maxima in sequences of random variables. Theory of Probability and its Applications, 56 (4):590–601, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zeifman:2017:TBA

[ZKKS17] A. I. Zeifman, A. V. Korotysheva, V. Yu. Korolev, and Ya. A. Satin. Truncation bounds for approximations of inhomogeneous continuous-time Markov chains. Theory of Probability and its Applications, 61(3):513–520, ???? 2017. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhitlukhin:2013:CHT

[ZM13] M. V. Zhitlukhin and A. A. Muravlev. On Chernoff’s hypotheses testing problem for the drift of a Brownian motion. Theory of Probability and its Applications, 57(4):708–717, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zuoxiang:2011:ASL

[ZN11] P. Zuoxiang and S. Nadarajah. Almost sure limit theorems for Gaussian sequences. Theory of Probability and its Applications, 55(2):361–367, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/
1/tprbau/v55/i2/p361_s1.

Zorich:2015:MGF

[Zor15] V. A. Zorich. Multidimensional geometry, functions of very many variables, and probability. Theory of Probability and its Applications, 59(3):481–493, ???? 2015. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zorich:2018:GP

[Zor18] V. A. Zorich. Geometry and probability. Theory of Probability and its Applications, 62 (2):236–251, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhdanov:2018:HEG

[ZP18] A. I. Zhdanov and V. I. Piterbarg. High extremes of Gaussian chaos processes: a discrete time approximation approach. Theory of Probability and its Applications, 63(1):1–21, ???? 2018. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhitlukhin:2013:BDP

[ZS13a] M. V. Zhitlukhin and A. N. Shiryaev. Bayesian disorder problems on filtered probability spaces. Theory of Probability and its Applications, 57(3):497–511, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zubkov:2013:CPU

[ZS13b] A. M. Zubkov and A. A. Serov. A complete proof of universal inequalities for the distribution function of the binomial law. Theory of Probability and its Applications, 57(3):539–544, ???? 2013. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhitlukhin:2014:OSP

[ZS14] M. V. Zhitlukhin and A. N. Shiryaev. Optimal stopping problems for a Brownian motion with disorder on a segment. Theory of Probability and its Applications, 58(1):164–171, ???? 2014. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zou:2019:RPD

[ZX19] W. Zou and J. H. Xie. A risk process with delayed claims and constant dividend barrier. Theory of Probability and its Applications, 64(1):103–123, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).